### v2

Abstract
The Earth Mover’s Distance is a popular similarity measure in several branches of computer science. It measures the minimum total edge length of a perfect matching between two point sets. The Earth Mover’s Distance under Translation ($\mathrm{EMDuT}$) is a translation-invariant version thereof. It minimizes the Earth Mover’s Distance over all translations of one point set.
For $\mathrm{EMDuT}$ in $\mathbb{R}^{1}$, we present an $\widetilde{\mathcal{O}}(n^{2})$-time algorithm. We also show that this algorithm is nearly optimal by presenting a matching conditional lower bound based on the Orthogonal Vectors Hypothesis. For $\mathrm{EMDuT}$ in $\mathbb{R}^{d}$, we present an $\widetilde{\mathcal{O}}(n^{2d+2})$-time algorithm for the $L_{1}$ and $L_{\infty}$ metric. We show that this dependence on $d$ is asymptotically tight, as an $n^{o(d)}$-time algorithm for $L_{1}$ or $L_{\infty}$ would contradict the Exponential Time Hypothesis (ETH). Prior to our work, only approximation algorithms were known for these problems.
[Uncaptioned image][Uncaptioned image]
1 Introduction
Earth Mover’s Distance (EMD).
$\mathrm{EMD}$, also known as geometric transportation or geometric bipartite matching, is a widely studied distance measure (see, e.g., [Indyk07, AndoniIK08, AndoniBIW09, AndoniNOY14, Rohatgi19, KhesinNP20, FoxL22, FoxL23, AgarwalCRX22, AgarwalRSS22]) that has received significant interest in computer vision, starting with the work of [RubnerTG00]. Depending on the precise formulation, $\mathrm{EMD}$ is a distance measure on point sets, distributions, or functions. In this paper, we study the following formulation of $\mathrm{EMD}$ as measuring the distance from a set of blue points $B$ to a set of red points $R$:
$\mathrm{EMD}_{p}(B,R)=\min_{\text{injective }\phi\colon B\to R}\sum_{b\in B}\|b-\phi(b)\|_{p}.$
Here, the minimization goes over all injective functions from $B$ to $R$, i.e., $\phi$ encodes a perfect matching of the points in $B$ to points in $R$, and the cost of a matching is the total length of all matching edges, with respect to the $L_{p}$ metric, $1\leq p\leq\infty$. When the value of $p$ is irrelevant, we may drop the subscript $p$.
The $\mathrm{EMD}_{p}$ problem is to compute the value $\mathrm{EMD}_{p}(B,R)$ for given sets $B,R\subseteq\mathbb{R}^{d}$ of sizes $|B|\leq|R|=n$. This general problem is sometimes called the asymmetric $\mathrm{EMD}$. The symmetric $\mathrm{EMD}$ is the special case with the additional restriction $|B|=|R|$. Intuitively, the asymmetric $\mathrm{EMD}$ asks whether $B$ is similar to some subset of $R$, while the symmetric variant compares the full sets $B$ and $R$. In this paper, we assume the dimension $d$ to be constant.
We briefly discuss algorithms for $\mathrm{EMD}$. Note that $\mathrm{EMD}$ can be formulated as a mincost matching problem on a bipartite graph with vertices $R\cup B$, where edge lengths are equal to the point-to-point distances. This graph has $|R|\cdot|B|=\mathcal{O}(n^{2})$ edges and solving bipartite mincost matching by the Hungarian method yields an exact algorithm for $\mathrm{EMD}$ with running time $\mathcal{O}(n^{3})$. Alternatively, by combining geometric spanners with recent advancements in (approximate) mincost flow solvers, one can obtain fast approximation algorithms for $\mathrm{EMD}$. For instance, symmetric $\mathrm{EMD}$ in $L_{2}$ metric can be solved in time $n(\log(n)/\varepsilon)^{\mathcal{O}(d)}$ [KhesinNP20]. See also [Indyk07, AndoniNOY14, FoxL22, FoxL23, AgarwalCRX22, AgarwalRSS22] for more approximation algorithms. Conditional lower bounds are also known, but only when the dimension is super-constant [Rohatgi19].
Earth Mover’s Distance under Translation (EMDuT).
We study a variant of $\mathrm{EMD}$ that is invariant under translations, and thus compares shapes of point sets, ignoring their absolute positions:
$\mathrm{EMDuT}_{p}(B,R)=\min_{\tau\in\mathbb{R}^{d}}\mathrm{EMD}_{p}(B+\tau,R).$
Here, $B+\tau=\{b+\tau\mid b\in B\}$ is the translated point set. See Figure 1.1 for an illustration of this distance measure. Again, we call asymmetric $\mathrm{EMDuT}_{p}$ the problem of computing $\mathrm{EMDuT}_{p}(B,R)$ for given sets $B,R$ of size $|B|\leq|R|=n$, and the symmetric variant comes with the additional restriction $|B|=|R|$. This measure was introduced by Cohen and Guibas [CohenG99], who presented heuristics as well as an exact algorithm with respect to the squared Euclidean distance. Later, Klein and Veltkamp [KleinV05] designed a 2-approximation algorithm for symmetric $\mathrm{EMDuT}_{p}$ running in asymptotically the same time as any $\mathrm{EMD}$ algorithm. Cabello, Giannopoulos, Knauer, and Rote [CabelloGKR08] designed $(1+\varepsilon)$-approximation algorithms for $\mathrm{EMDuT}_{2}$ in the plane, running in time $\widetilde{\mathcal{O}}(n^{4}/\varepsilon^{4})$ for the asymmetric variant and $\widetilde{\mathcal{O}}(n^{3/2}/\varepsilon^{7/2})$ for the symmetric variant. [^1] Here and throughout the paper we use $\widetilde{\mathcal{O}}$ notation to ignore logarithmic factors, i.e., $\widetilde{\mathcal{O}}(T)=\bigcup_{c\geq 0}\mathcal{O}(T(\log T)^{c})$. Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] proposed algorithms to solve the symmetric $\mathrm{EMDuT}_{1}$ and symmetric $\mathrm{EMDuT}_{\infty}$ problems in the plane, that run in $\mathcal{O}(n^{6}\log^{3}n)$ time. We remark that most of these works also study variants of $\mathrm{EMDuT}$ under more general transformations than translations, but in this paper we focus on translations.
We are not aware of any other research on $\mathrm{EMDuT}$, which is surprising, since translation-invariant distance measures are well motivated, and the analogous Hausdorff distance under translation [HuttenlocherK90, Rote91, HuttenlocherRK92, AgarwalHSW10, KnauerS11, KnauerKS11, BringmannN21, Chan23] and Fréchet distance under translation [AltKW01, MosigC05, JiangXZ08, AvrahamKS15, BringmannKN20, FiltserK20, BringmannKN21] have received considerably more attention.
Figure 1.1: Given a set of (solid) blue points $B$ and a set of red points $R$, our goal is to find a translation $\tau$ (shown in green) and a perfect matching from $B+\tau$ to $R$ (shown in black) that minimizes the total distance of matched pairs.
1.1 Our Results
We study $\mathrm{EMDuT}$ from the perspective of fine-grained complexity. We design new algorithms and prove conditional lower bounds over $\mathbb{R}^{1}$, as well as for $L_{1}$ and $L_{\infty}$ over $\mathbb{R}^{d}$.
EMDuT in 1D.
Over $\mathbb{R}^{1}$ all $L_{p}$ metrics are equal. We present the following new algorithms.
Theorem 1.1 (1D Algorithms).
(Symmetric:) Given sets $B,R\subseteq\mathbb{R}$ of size $n=|B|=|R|$, $\mathrm{EMDuT}(B,R)$ can be computed in time $\mathcal{O}(n\log n)$. (Asymmetric:) Given sets $B,R\subseteq\mathbb{R}$ of size $m=|B|\leq n=|R|$, $\mathrm{EMDuT}(B,R)$ can be computed in time $\mathcal{O}(mn(\log n+\log^{2}m))$.
Note that for $m=\Omega(n)$, for the asymmetric variant we obtain near-quadratic time $\widetilde{\mathcal{O}}(n^{2})$, while for the symmetric variant we obtain near-linear time $\widetilde{\mathcal{O}}(n)$. We fully explain this gap, by proving a matching conditional lower bound showing that no algorithm solves the asymmetric variant in strongly subquadratic time $\mathcal{O}(n^{2-\delta})$ for any $\delta>0$, for $m=\Omega(n)$. In fact, we present a stronger lower bound that even rules out fast approximation algorithms, not only fast exact algorithms. Our lower bound assumes the Orthogonal Vectors Hypothesis (OVH), a widely-accepted conjecture from fine-grained complexity theory; for a definition see Section 4.
Theorem 1.2 (1D Lower Bound).
Assuming OVH, for any constant $\delta>0$ there is no algorithm that, given $\varepsilon\in(0,1)$ and sets $B,R\subseteq\mathbb{R}$ of size $n=|R|\geq|B|=\Omega(n)$, computes a $(1+\varepsilon)$-approximation of $\mathrm{EMDuT}(B,R)$ in time $\mathcal{O}(n^{2-\delta}/\varepsilon^{o(1)})$.
As a corollary, the same conditional lower bound holds for $\mathrm{EMDuT}_{p}$ over $\mathbb{R}^{d}$, for any $d\geq 1$ and $1\leq p\leq\infty$, since subsets of $\mathbb{R}$ can be embedded into $\mathbb{R}^{d}$ for any dimension $d$ and any $L_{p}$ metric.
Let us give a brief overview of these results. In the symmetric setting, we establish that $f(\tau)\coloneq\mathrm{EMD}(B+\tau,R)$ is a unimodal function in $\tau$, i.e., it is first monotone decreasing and then monotone increasing, and thus its minimum can be found easily. In contrast, in the asymmetric setting the function $f(\tau)$ can have up to $\Theta(n^{2})$ disconnected global minima. Intuitively, our lower bound shows that any algorithm needs to consider each one of these global near-minima, and therefore the running time must be quadratic in order to determine which near-minimum is the actual global minimum. To obtain our algorithm in the asymmetric setting, we use a sweep algorithm with an intricate event handling data structure.
EMDuT for $L_{1}$ and $L_{\infty}$ metric in higher dimensions.
We extend the work of Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] for point sets in $\mathbb{R}^{d}$, leading to the following algorithms.
Theorem 1.3 (Algorithms for $L_{1}$ and $L_{\infty}$ metric, Asymmetric).
Given sets $B,R\subseteq\mathbb{R}^{d}$ of size $m=|B|\leq n=|R|$, $\mathrm{EMDuT}_{1}(B,R)$ and $\mathrm{EMDuT}_{\infty}(B,R)$ can be computed in $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time.
We explain that such a dependence on the dimension is unavoidable, by establishing a more coarse-grained lower bound compared to our 1D results: We show that no algorithm can solve the problem in time $n^{o(d)}$. In fact, we present a stronger lower bound that even rules out fast approximation algorithms. Our lower bound assumes the Exponential Time Hypothesis (ETH) [ImpagliazzoP01], which is a well-established conjecture from fine-grained complexity theory.
Theorem 1.4 (Lower Bound for $L_{1}$ and $L_{\infty}$ metric, Symmetric).
Assuming ETH, there is no algorithm that, given $\varepsilon\in(0,1)$ and sets $B,R\subseteq\mathbb{R}^{d}$ of size $n=|B|=|R|$, computes a $(1+\varepsilon)$-approximation of $\mathrm{EMDuT}_{1}(B,R)$ in time $(\frac{n}{\varepsilon})^{o(d)}$. The same holds for $\mathrm{EMDuT}_{\infty}(B,R)$.
Note that our lower bound pertains to the symmetric setting, while our algorithm addresses the more general asymmetric setting. Hence, these results together cover both the symmetric and the asymmetric setting.
Let us give a brief overview of these results. For the algorithm, we establish an arrangement of complexity $\mathcal{O}(m^{d}n^{d})$ such that the optimal translation $\tau$ is attained at one of the vertices within this arrangement. Our algorithm is obtained by computing the $\mathrm{EMD}$ at each vertex. The lower bound is proven via a reduction from the $k$-Clique problem. In our construction, each coordinate of the translation $\tau$ chooses one vertex from a given $k$-Clique instance. We design gadgets that verify that every pair of selected nodes indeed forms an edge.
1.2 Open Problems
EMDuT in 1D.
Over $\mathbb{R}^{1}$, we leave open whether there are fast approximation algorithms: Can a constant-factor approximation be computed in time $\mathcal{O}(n^{2-\delta})$ for some constant $\delta>0$? Or even in time $\widetilde{\mathcal{O}}(n)$? Can a $(1+\varepsilon)$-approximation be computed in time $\widetilde{\mathcal{O}}(n^{2-\delta}/\textup{poly}(\varepsilon))$ for some constant $\delta>0$ (independent of $n$ and $\varepsilon$)? Or even in time $\widetilde{\mathcal{O}}(n/\textup{poly}(\varepsilon))$?
EMDuT for $L_{1}$ and $L_{\infty}$ metric in higher dimensions.
For the $L_{1}$ and $L_{\infty}$ metric in dimension $d\geq 2$ we leave open to determine the optimal constant $c>0$ such that the problem can be solved in time $n^{c\cdot d+o(d)}$.
EMDuT for $L_{2}$ metric in higher dimensions.
The $L_{2}$ metric is the most natural measure in geometric settings, making $\mathrm{EMDuT}_{2}$ a well motivated problem. The most pressing open problem is to determine the complexity of the $\mathrm{EMDuT}_{2}$ problem in any dimension $d\geq 2$.
We first observe that the $\mathrm{EMDuT}_{2}$ problem cannot be solved exactly (on the Real RAM model of computation). Indeed, on the Real RAM supporting only the usual arithmetic operations ($+,-,\cdot,/$), if the input numbers are rational then all output numbers are rational. If the machine further supports square roots (or other integral roots), if the input numbers are rational then the output numbers are algebraic. For the Geometric Median problem there are instances with rational input coordinates where the coordinates of the geometric median are not algebraic; this even holds in two dimensions [Bajaj88]. Therefore, Geometric Median cannot be solved exactly on the Real RAM. Finally, note that Geometric Median is a special case of $\mathrm{EMDuT}_{2}$, as for any point set $R\subset\mathbb{R}^{d}$ of size $n$, if $B$ consists of $n$ copies of the point $(0,\ldots,0)$, then $\mathrm{EMDuT}_{2}(B,R)$ is the (cost of the) Geometric Median of $R$. Therefore, also $\mathrm{EMDuT}_{2}$ cannot be solved exactly on the Real RAM.
We therefore need to relax the goal and ask for an approximation algorithm. Geometric Median has a very fast $(1+\varepsilon)$-approximation algorithm running in time $\mathcal{O}(nd\log^{3}(1/\varepsilon))$ [CohenLMPS16], so the reduction from Geometric Median to $\mathrm{EMDuT}_{2}$ does not rule out very fast approximation algorithms for $\mathrm{EMDuT}_{2}$.
This is in stark contrast to what we know about the $\mathrm{EMDuT}_{2}$ problem, as almost all of our techniques in this paper completely fail for this problem. We neither obtain an algorithm running in time $n^{\mathcal{O}(d)}$, nor can we prove a lower bound ruling out time $n^{o(d)}$. On the lower bound side, all we know is the lower bound from 1D, ruling out $(1+\varepsilon)$-approximation algorithms running in time $\mathcal{O}(n^{2-\delta}/\varepsilon^{o(1)})$ for any constant $\delta>0$. On the algorithms side, one can observe that after fixing the matching from $B$ to $R$, the problem of finding the optimal translation $\tau$ for this matching is the Geometric Median problem and thus has a $(1+\varepsilon)$-approximation algorithm running in time $\mathcal{O}(nd\log^{3}(1/\varepsilon))$. By trying out all $n^{\mathcal{O}(n)}$ possible matchings, one can obtain a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{2}$ running in time $n^{\mathcal{O}(n)}\log^{3}(1/\varepsilon)$ for any constant $d$. We pose as an open problem to close this huge gap between the quadratic lower and exponential upper bound (for $(1+\varepsilon)$-approximation algorithms with a $1/\varepsilon^{o(1)}$ dependency on $\varepsilon$ in the running time).
2 Preliminaries
We use $[n]$ to denote $\{1,\ldots,n\}$. All logarithms are base $2$. For every $x\in\mathbb{R}$ we let $\lfloor x\rceil\in\mathbb{Z}$ be the unique integer such that $x-\lfloor x\rceil\in(-1/2,1/2]$. Consider a set of blue points $B\subseteq\mathbb{R}^{d}$ and a set of red points $R\subseteq\mathbb{R}^{d}$. Fix an $L_{p}$ norm, for any $1\leq p\leq\infty$. Denote by $\Phi$ the set of all injective functions $\phi\colon B\to R$, i.e., $\Phi$ is the set of all perfect matchings from $B$ to $R$. For any matching $\phi\in\Phi$ and any translation $\tau\in\mathbb{R}^{d}$ we define the cost
$\mathcal{D}_{B,R,p}(\phi,\tau)=\sum_{b\in B}\|b+\tau-\phi(b)\|_{p}.$
We will ignore the subscript $p$ when it is clear from the context. Note that we can express $\mathrm{EMD}$ and $\mathrm{EMDuT}$ in terms of this cost function as
$\mathrm{EMD}_{p}(B,R)=\min_{\phi\in\Phi}\mathcal{D}_{B,R,p}(\phi,(0,\ldots,0))\,\text{ and }\,\mathrm{EMDuT}_{p}(B,R)=\min_{\phi\in\Phi}\min_{\tau\in\mathbb{R}^{d}}\mathcal{D}_{B,R,p}(\phi,\tau).$
3 Algorithm in One Dimension
We first consider computing $\mathrm{EMDuT}_{p}(B,R)$ for two point sets $B,R$ in $\mathbb{R}^{1}$. For ease of presentation, assume that $R$ and $B$ are indeed sets, and thus there are no duplicate points. We can handle the case of duplicate points by symbolic perturbation. Observe, that the distance between a pair of points $b,r$ in any $L_{p}$ metric is simply $\|b-r\|_{p}=\|b-r\|_{1}=|b-r|$. In Section 3.1, we describe a very simple $\mathcal{O}(n\log n)$ time algorithm to compute $\mathrm{EMDuT}_{p}(B,R)$ (as well as an optimal matching $\phi^{*}$ and translation $\tau^{*}$ that realize this distance) when $B$ and $R$ both contain exactly $n$ points. In Section 3.2, we consider the much more challenging case where $|B|=m$ and $|R|=n$ differ. For this case we develop an $\mathcal{O}(nm(\log n+\log^{2}m))$ time algorithm to compute $\mathrm{EMDuT}_{p}(B,R)$.
A matching $\phi$ is said to be monotonically increasing if and only if for every pair of blue points $b^{\prime}<b$ we also have $\phi(b^{\prime})<\phi(b)$. We show the following crucial property.
Lemma 3.1.
For any $B,R\subset\mathbb{R}$ there is an optimal matching $\phi$ that is monotonically increasing.
Proof.
We say that $(b,b^{\prime})\in B\times B$ forms a crossing in a matching $\phi$ if $b>b^{\prime}$ and $\phi(b)<\phi(b^{\prime})$. Let $\phi$ be an optimal matching of $B,R\subseteq\mathbb{R}$ with the minimal number of crossings. If $\phi$ does not have any crossing, it is monotonically increasing. Hence, for the sake of contradiction assume that $(b,b^{\prime})$ is a crossing in $\phi$. Let $r=\phi(b^{\prime})$ and $r^{\prime}=\phi(b)$ and consider a matching $\phi^{\prime}$ that has $\phi^{\prime}(b)=r$ and $\phi^{\prime}(b^{\prime})=r^{\prime}$, and $\phi^{\prime}(x)=\phi(x)$ for every $x\in B\setminus\{b,b^{\prime}\}$.
We will show that $\mathcal{D}_{B,R}(\phi,0)\geq\mathcal{D}_{B,R}(\phi^{\prime},0)$. Combined with the fact that $\phi^{\prime}$ has less crossings than $\phi$, this yields a contradiction to the choice of $\phi$ as the optimal matching with the minimal number of crossings.
Note that $\mathcal{D}_{B,R}(\phi,0)\geq\mathcal{D}_{B,R}(\phi^{\prime},0)$ is equivalent to
$|b-r^{\prime}|+|b^{\prime}-r|\geq|b-r|+|b^{\prime}-r^{\prime}|.$ (3.1)
Since $b^{\prime}<b$ and $r^{\prime}<r$, inequality (3.1) follows from the fact below (by setting $x=b^{\prime}-r$, $\alpha=b-b^{\prime}$, and $\beta=r-r^{\prime}$). ∎
Fact 3.2.
For every $x\in\mathbb{R}$ and $\alpha,\beta>0$ it holds that $|x|+|x+\alpha+\beta|\geq|x+\alpha|+|x+\beta|$.
3.1 Symmetric Case
In the symmetric case ($|R|=|B|$), Lemma 3.1 uniquely defines an optimal matching. Let $B=\{b_{1},\ldots,b_{n}\}$ and $R=\{r_{1},\ldots,r_{n}\}$ be the points in increasing order. Now, the optimal translation $\tau^{*}$ is the value for $\tau$ that minimizes $\mathcal{D}_{B,R}(\phi,\tau)=\sum_{i=1}^{n}|b_{i}-r_{i}+\tau|$. Thus, it corresponds to the median of $b_{1}-r_{1},\ldots,b_{n}-r_{n}$, which we can compute in $\mathcal{O}(n\log n)$ time.
Theorem 3.3.
We can compute $\mathrm{EMDuT}(R,B)$ in 1D in $\mathcal{O}(n\log n)$ time when $|R|=|B|$.
3.2 Asymmetric Case
In this section, we present an $\mathcal{O}(mn(\log n+\log^{2}m))$ time algorithm to compute $\mathrm{EMDuT}(B,R)$, for the case that $m\leq n$. Consider the cost $f(\tau)=\min_{\phi\in\Phi}\mathcal{D}_{B,R}(\phi,\tau)$ as a function of $\tau$. The minimum of this function is $\mathrm{EMDuT}(B,R)$. The main idea is then to sweep over the domain of $f$, increasing $\tau$ from -$\infty$ to $\infty$, while maintaining (a representation of) $f$ and a matching $\phi$ that realizes cost $f(\tau)=\mathcal{D}_{B,R}(\phi,\tau)$. We also maintain the best translation $\tau^{*}\leq\tau$ (i.e. with minimal cost) among the translations considered so far (and if there are multiple such translations, the smallest one), so at the end of our sweep, $\tau^{*}$ is thus an optimal translation.
Properties of $f$.
By Lemma 3.1, for any $\tau$, there exists an optimal monotonically increasing matching between $B+\tau$ and $R$. So, we restrict our attention to such monotonically increasing matchings. Observe that any such matching $\phi$ corresponds to a partition of $B$ into runs, i.e. maximal subsequences of consecutive points, $B_{1},\ldots,B_{z}$, so that the points $b_{t-k},\ldots,b_{t}$ in a run $B_{i}$ are matched to consecutive red points $r_{u-k},\ldots,r_{u}$, for some $r_{u}=\phi(b_{t})$. Moreover, for any such a matching $\phi$, the function $\mathcal{D}_{B,R}(\phi,\tau)$ is piecewise linear in $\tau$, and each breakpoint is a translation $\tau$ for which there is a pair $(b,r)\in B\times R$ with $b+\tau=r$. It then follows that $f(\tau)$ is also piecewise linear in $\tau$. Furthermore, the breakpoints of $f$ are of two types. A type (i) breakpoint is a translation such that there is a pair $(b,r)\in B\times R$ with $b+\tau=r$, and a type (ii) breakpoint if there are two different matchings $\phi,\phi^{\prime}$ that both realize the same minimum cost $\mathcal{D}_{B,R}(\phi,\tau)=\mathcal{D}_{B,R}(\phi^{\prime},\tau)$. We show the following key lemma, which lets us characterize the breakpoints of type (ii) more precisely.
Figure 3.1: Schematic representation of the graph $G=\phi\oplus\phi^{\prime}$ used in the proof of Lemma˜3.4. Each edge exists if and only if exactly one edge from either $\phi$ or $\phi^{\prime}$ is present. Green edges arise from the matching $\phi^{\prime}$, while yellow edges arise from the matching $\phi$.
Lemma 3.4.
Let $\phi$ be an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau,R)$, and let $\phi^{\prime}$ be an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau^{\prime},R)$ for some $\tau^{\prime}>\tau$. Then, $\phi^{\prime}(b)\geq\phi(b)$ for all $b\in B$.
Proof.
We assume that all the points are distinct, as otherwise we can perturb them infinitesimally to resolve ties. Consider a bipartite graph $G$ defined as follows: the vertices of $G$ are $B\uplus R$, and we add an edge $(b,r)\in B\times R$ to $G$ if exactly one of the following conditions holds: (i) $\phi(b)=r$ and $\phi^{\prime}(b)\neq r$, or (ii) $\phi^{\prime}(b)=r$ and $\phi(b)\neq r$. The graph $G$ can be thought of as the exclusive-or of the matchings $\phi$ and $\phi^{\prime}$, see Figure˜3.1. We will now demonstrate that the connected components of this graph are paths. Then, considering that the matchings are monotone, it follows that the edges of these paths are non-crossing. This implies that consecutive red vertices on these paths are monotone. The lemma statement then easily follows.
Let $C$ be any connected component of $G$ that consists of more than one vertex.
Claim 3.5.
$C$ is a path.
Proof.
Observe that the maximum degree of graph $G$ is $2$, so the connected components of $G$ consist of cycles and paths. Assume that $C$ is a cycle. In that case, however, there exists a pair of edges from either $\phi$ or $\phi^{\prime}$ that intersect (see Figure˜3.2). This contradicts the assumption about the monotonicity of both $\phi$ and $\phi^{\prime}$. Hence, $C$ is not a cycle, and the proof of the claim follows. ∎
Figure 3.2: Case when connected component of $G$ is a cycle.
Now we know that $C$ is a path. Let $\{p_{1},\ldots,p_{\ell}\}=V(C)$ be the consecutive vertices on the path $C$. We have the freedom to select the order of endpoints; hence, without loss of generality, assume that $p_{1}$ is on the left of $p_{\ell}$, i.e., $p_{1}<p_{\ell}$.
Claim 3.6.
Vertices $p_{1}$ and $p_{\ell}$ are in $R$.
Proof.
Every vertex $b\in B$ has degree $2$ in $G$ unless $b$ is an isolated vertex. However, both vertices $p_{1}$ and $p_{\ell}$ are endpoints of a path, which means that $p_{1}$ and $p_{\ell}$ have degree $1$ in $G$. This means that $p_{1},p_{\ell}\notin B$. ∎
Since $G$ is a bipartite graph, this means that $\ell$ is odd. We assumed that $|V(C)|>1$, so $\ell\geq 3$. Observe that if $i$ is odd, then $p_{i}\in R$, and if $i$ is even, then $p_{i}\in B$. Now, we show that red and blue vertices in $C$ are monotone:
Claim 3.7.
For every $i\in[\ell-2]$, it holds that $p_{i}<p_{i+2}$.
Proof.
When $\ell=3$, the claim holds because we have assumed $p_{1}<p_{3}$. Hence, we can assume that $\ell\geq 5$ (since $\ell$ is odd). For the sake of contradiction, assume that $p_{i}>p_{i+2}$ for some $i$. Let $t_{i}\coloneqq p_{i+2}-p_{i}$ for $i\in[\ell-2]$. This means that there exists $i\in[\ell-2]$ with $t_{i}<0$. Moreover, we have $p_{\ell}>p_{1}$, which means that $\sum_{\text{odd }i}t_{i}>0$. Therefore, there also exists $j\in[\ell-2]$ with $t_{j}>0$.
In particular, there exists an index $k\in[\ell-2]$ such that $t_{k}\cdot t_{k+1}<0$. This means that either (a) $p_{k}>p_{k+2}$ and $p_{k+1}<p_{k+3}$, or (b) $p_{k}<p_{k+2}$ and $p_{k+1}>p_{k+3}$. In both of these cases, the intervals $(\min\{p_{k},p_{k+1}\},\max\{p_{k},p_{k+1}\})$ and $(\min\{p_{k+2},p_{k+3}\},\max\{p_{k+2},p_{k+3}\})$ intersect (see Figure˜3.3), which contradicts the assumption about the monotonicity of $\phi$ and $\phi^{\prime}$. ∎
Figure 3.3: Two cases of Claim˜3.7 in which a crossing occurs.
Now, we continue the proof of Lemma˜3.4. For the sake of contradiction, assume that $\phi^{\prime}(b)<\phi(b)$ for some $b\in B$. By Claim˜3.7, there exists a connected component $C$ of $G$ such that $\phi^{\prime}(b)<\phi(b)$ for every $b\in C\cap B$. Fix any such component $C$ and for every $b\in B$ let:
$\psi(b)\coloneq\begin{cases}\phi^{\prime}(b)&\text{ if }b\in C\cap B,\\ \phi(b)&\text{ otherwise.}\end{cases}$ and $\psi^{\prime}(b)\coloneq\begin{cases}\phi(b)&\text{ if }b\in C\cap B,\\ \phi^{\prime}(b)&\text{ otherwise.}\end{cases}$
Notice that $\psi$ and $\psi^{\prime}$ are both matchings of $\mathrm{EMD}_{p}(B+\tau,R)$ and $\mathrm{EMD}_{p}(B+\tau^{\prime},R)$ respectively. Now, recall that the cost of each matching is:
$\mathcal{D}_{B,R}(\phi,\tau)=$ $\sum_{b\in C\cap B}\lvert b+\tau-\phi(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau-\phi(b)\rvert$,
$\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})=$ $\sum_{b\in C\cap B}\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert$,
$\mathcal{D}_{B,R}(\psi,\tau)=$ $\sum_{b\in C\cap B}\lvert b+\tau-\phi^{\prime}(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau-\phi(b)\rvert,\text{ and }$
$\mathcal{D}_{B,R}(\psi^{\prime},\tau^{\prime})=$ $\sum_{b\in C\cap B}\lvert b+\tau^{\prime}-\phi(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert$.
Since $\phi$ and $\phi^{\prime}$ are optimal matchings for $\tau$ and $\tau^{\prime}$ respectively, we have:
$\mathcal{D}_{B,R}(\phi,\tau)\leq\mathcal{D}_{B,R}(\psi,\tau)$ and $\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})\leq\mathcal{D}_{B,R}(\psi^{\prime},\tau^{\prime}).$ (3.2)
We say that $b\in B$ is a crossing if $\phi^{\prime}(b)<\phi(b)$. Notice that every $b\in C\cap B$ is a crossing. Next, we classify crossings into types based on the order of the points:
  • Type BBRR: $b+\tau<b+\tau^{\prime}\leq\phi^{\prime}(b)<\phi(b)$,
  • Type RRBB: $\phi^{\prime}(b)<\phi(b)\leq b+\tau<b+\tau^{\prime}$,
  • Type RBBR: $\phi^{\prime}(b)\leq b+\tau<b+\tau^{\prime}\leq\phi(b)$,
  • Type BRBR: $b+\tau\leq\phi^{\prime}(b)<b+\tau^{\prime}\leq\phi(b)$,
  • Type RBRB: $\phi^{\prime}(b)\leq b+\tau<\phi(b)\leq b+\tau^{\prime}$,
  • Type BRRB: $b+\tau\leq\phi^{\prime}(b)<\phi(b)\leq b+\tau^{\prime}$.
Figure 3.4: Illustration of crossing types. The left figure shows a BBRR crossing and the right figure shows a RBBR crossing.
Note that for any crossing $b$ of type BBRR or RRBB it holds that:
$|\phi(b)-b-\tau|+|\phi^{\prime}(b)-b-\tau^{\prime}|=|\psi(b)-b-\tau|+|\psi^{\prime}(b)-b-\tau^{\prime}|.$
Moreover, for any crossing $b$ of type RBBR, BRBR, RBRB or BRRB, denoting by $x_{1}\leq x_{2}<x_{3}\leq x_{4}$ the numbers $b+\tau,b+\tau^{\prime},\phi(b)$ and $\phi^{\prime}(b)$ in sorted order, we have:
$|\phi(b)-b-\tau|+|\phi^{\prime}(b)-b-\tau^{\prime}|$ $=x_{3}+x_{4}-x_{1}-x_{2},$
$|\psi(b)-b-\tau|+|\psi^{\prime}(b)-b-\tau^{\prime}|$ $=x_{2}+x_{4}-x_{1}-x_{3}.$
Since $x_{3}>x_{2}$, it follows that for any crossing $b$ of type RBBR, BRBR, RBRB or BRRB we have:
$|\phi(b)-b-\tau|+|\phi^{\prime}(b)-b-\tau^{\prime}|>|\psi(b)-b-\tau|+|\psi^{\prime}(b)-b-\tau^{\prime}|.$
By summing up these inequalities over all $b\in C\cap B$, if at least one crossing in $C$ has type RBBR, BRBR, RBRB or BRRB we obtain
$\mathcal{D}_{B,R}(\phi,\tau)+\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})>\mathcal{D}_{B,R}(\psi,\tau)+\mathcal{D}_{B,R}(\psi^{\prime},\tau^{\prime}),$
contradicting inequalities (3.2).
It remains to consider the case that all crossings in $C$ have type BBRR or RRBB. Note that if the leftmost crossing $b=\min(B\cap C)$ has type BBRR, then we can improve the matching $\phi$ by changing $\phi(b)$ to $\phi^{\prime}(b)$; this contradicts the assumption that $\phi$ is an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau^{\prime},R)$. Symmetrically, we obtain a contradiction if the rightmost crossing $b=\max(B\cap C)$ has type RRBB. Finally, note that since $C$ is a connected component and both $\phi$ and $\phi^{\prime}$ are monotone, if $C$ has a leftmost crossing of type RRBB and a rightmost crossing of type BBRR, then in between these two there must be at least one crossing of type RBBR, BRBR, RBRB or BRRB. As this case was handled in the previous paragraph, we finished the proof. ∎
We can now come back to breakpoints and runs of $\phi$.
Corollary 3.8.
A breakpoint $\tau$ of type (ii) corresponds to a pair of optimal monotonically increasing matchings $\phi,\phi^{\prime}$ for which for all points $b\in B$ we have $\phi(b)\leq\phi^{\prime}(b)$. Furthermore, consider a run $b_{s},\ldots,b_{t}$ of $\phi$ and a point $b_{i}$ with $i\in\{s,\ldots,t\}$. If $\phi(b_{i})<\phi^{\prime}(b_{i})$, then $\phi(b_{j})<\phi^{\prime}(b_{j})$ for all $j\in\{i,\ldots,t\}$.
Proof.
Recall that $\tau$ is a breakpoint of type (ii) if there are two distinct matchings $\phi$ and $\phi^{\prime}$ that both realize the optimal cost at $\tau$. Assume that $\phi$ is also optimal for a translation $\tau^{\prime}<\tau$ or that $\phi^{\prime}$ is also optimal for a translation $\tau^{\prime\prime}>\tau$. Now by Lemma 3.4 we have $\phi(b)\leq\phi^{\prime}(b)$ for every $b\in B$. Next, recall that $b_{s},\dots,b_{t}$ form a run if $\phi$ matches them to consecutive red points. Since $\phi(b_{i+1})$ is matched to the red point right after $\phi(b_{i})$, now $\phi^{\prime}(b_{i})>\phi(b_{i})$ implies in particular that $\phi^{\prime}(b_{i})\geq\phi(b_{i+1})$. With monotonicity this gives us that $\phi^{\prime}(b_{i+1})>\phi^{\prime}(b_{i})\geq\phi(b_{i+1})$. By induction, the same holds for the remaining $j\in\{i+2,\dots,t\}$. ∎
Lemma 3.9.
The function $f(\tau)$ is piecewise linear, and consists of $\mathcal{O}(nm)$ pieces.
Proof.
As argued above, $f$ is piecewise linear. What remains is to argue that there are $\mathcal{O}(nm)$ breakpoints. For every pair of points $(b_{i},r_{j})\in B\times R$ there is only one translation $\tau$ such that $b+\tau=r$, so clearly there are at most $\mathcal{O}(nm)$ breakpoints of type (i). At every breakpoint of type (ii), there is at least one blue point $b_{i}$ that was matched to $r_{j}$ and gets matched to some $r_{k}$ with $k>j$. This also happens at most once for every pair $b_{i},r_{j}$. Hence, the number of breakpoints of type (ii) is also $\mathcal{O}(nm)$. ∎
In our sweep line algorithm, we will maintain a current optimal matching $\phi$. At each breakpoint of type (i) we will have an event to update the cost function of the matching. Furthermore, it follows from Corollary 3.8 that when we sweep over a breakpoint of type (ii), we can decompose the changes to the matching using a series of atomic events. In each such atomic event there is some suffix $b_{j},\ldots,b_{t}$ of a run $b_{s},\ldots,b_{t}$ that $\phi$ currently matches to $r_{u-t+j},\dots,r_{u}$ that will become matched to $r_{u-t+j+1},\ldots,r_{u+1}$. As we argued in the proof of Lemma 3.9, the total number of such events is only $\mathcal{O}(nm)$. Next, we express how we can efficiently compute the next such atomic event, and handle it.
Consider a run $B_{i}=b_{s},\ldots,b_{t}$ induced by $\phi$ at time $\tau$. Our aim is to find the smallest $\tau^{\prime}\geq\tau$ at which there is an atomic type (ii) event involving a suffix $b_{j},\ldots,b_{t}$ of $B_{i}$. Hence, for a given suffix $b_{j},\ldots,b_{t}$, we wish to maintain when it starts being beneficial to match $b_{j},\ldots,b_{t}$ to $r_{u-t+j+1},\ldots,r_{u+1}$ rather than to $r_{u-t+j},\ldots,r_{u}$.
Let $\Delta^{\prime}_{j}$ represent the change in cost when we match $b=b_{j}$ to $r^{\prime}=r_{v+1}$ rather than to $r=r_{v}$, ignoring that $r_{v+1}$ may already be matched to some other blue point. We have that
$\Delta^{\prime}_{j}(\tau)=|b-r^{\prime}+\tau|-|b-r+\tau|=\begin{cases}r^{\prime}-r&\text{if }b+\tau\leq r,\\ r+r^{\prime}-2b-2\tau&\text{if }r<b+\tau<r^{\prime},\\ r-r^{\prime}&\text{if }b+\tau\geq r^{\prime}.\end{cases}$
Figure 3.5: Each point $b_{j}$ in a run $B_{i}=b_{s},\ldots,b_{t}$ defines a (piecewise)-linear function $\Delta^{\prime}_{j}$. Each suffix $b_{j},\ldots,b_{t}$ then defines a linear function $\Delta_{j}$, expressing the cost of switching from matching $\phi$ to $\phi^{\prime}$. The lower envelope $g_{i}$ of these functions then defines the first type (ii) event $\tau^{\prime}$ of run $B_{i}$.
Observe that this function is piecewise linear, and non-increasing. Moreover, the breakpoints coincide with type (i) breakpoints of $f$ at which $b+\tau$ coincides with a red point. Hence, in between any two consecutive events, we can consider $\Delta^{\prime}_{j}$ as a linear function. See Figure 3.5 for an illustration.
We can then express the cost of changing the matching for the entire suffix $b_{j},\ldots,b_{t}$ as $\Delta_{j}(\tau)=\sum_{k=j}^{t}\Delta^{\prime}_{k}(\tau)$. This function is again decreasing, piecewise linear, and has breakpoints that coincide with type (i) breakpoints of $f$. When $\Delta_{j}(\tau)$ becomes non-positive it becomes beneficial to match the suffix $b_{j},\ldots,b_{t}$ to $r_{u-j+1},\ldots,r_{u+1}$. Hence, the first such translation is given by a root of $\Delta_{j}(\tau)$. Note that there is at most one such root since $\Delta_{j}$ is decreasing.
It now follows that (if it exists) the root $\tau^{\prime}$ of the function $g_{i}(\tau)=\min_{j\in\{s,\ldots,t\}}\Delta_{j}(\tau)$ expresses the earliest time that there is a suffix $b_{j},\ldots,b_{t}$ for which it is beneficial to update the matching. As before, this function is decreasing and piecewise linear. Hence, we obtain:
Lemma 3.10.
Let $\tau_{1},\tau^{\prime}]\ni\tau$ be a maximal interval on which $f(\tau)$ is linear, let $\tau^{\prime}$ be a type (ii) breakpoint, and let $\phi$ be an optimal matching for $\tau$. Then there is a run $B_{i}$ induced by $\phi$, and $\tau^{\prime}$ is a root of the function $g_{i}(\tau)$.
Representing the lower envelope $g_{i}$.
At any moment of our sweep, we maintain a single piece of $g_{i}$. Hence, this piece is the lower envelope of a set of linear functions $\Delta_{s},\ldots,\Delta_{t}$. We will maintain this lower envelope using an adapted version of the data structure by Overmars and van Leeuwen [overmars1981]. Ideally, we would maintain the lower envelope of $\Delta_{s},\ldots,\Delta_{t}$ directly. However, reassigning a single blue point $b_{j}$ in the matching $\phi$, may cause many functions $\Delta_{k}$ to change. So, we implicitly represent each function $\Delta_{j}$ as a sum of $\Delta^{\prime}_{k}$ functions.
Lemma 3.11.
Let $B_{i}$ be a run of size $k$. Using $\mathcal{O}(k\log k)$ space, we can represent the current piece of the lower envelope $g_{i}$ such that we can find the root of (this piece of) $g_{i}$ in $\mathcal{O}(\log k)$ time, and insert or remove any point in $B_{i}$ in $\mathcal{O}(\log^{2}k)$ time.
Proof.
We will maintain this lower envelope using (a slightly adapted version of) the data structure by Overmars and van Leeuwen [overmars1981]. They present a data structure to store a lower envelope of $k$ lines while allowing queries such as line intersections in $\mathcal{O}(\log k)$ time, as well as insertions and deletions in $\mathcal{O}(\log^{2}k)$ time. We could insert the lines (representing) $\Delta_{s},\ldots,\Delta_{t}$ into this data structure so we can efficiently find the root of $g_{i}$. However, an update to $\phi$ causes an update to a function $\Delta^{\prime}_{k}$, which in turn could cause changes in many functions $\Delta_{j}$, and could require significant changes to $g_{i}$. We therefore extend the data structure of Overmars and van Leeuwen to support such updates efficiently.
The Overmars and van Leeuwen data structure.
The data structure by Overmars and van Leeuwen is a balanced binary tree, whose leaves store the lines $\Delta_{s},\ldots,\Delta_{t}$ in the decreasing slope order. For each node $v$, let $\Delta^{v}$ denote the (ordered) set of lines stored in the leaves below $v$, and let $E^{v}$ denote the ordered set of lines that appear in left-to-right order on the lower envelope $\mathcal{E}^{v}$ of $\Delta^{v}$. Overmars and van Leeuwen observe that (i) this set of lines $E^{v}$ defining the lower envelope $\mathcal{E}^{v}$ actually forms an ordered subset of the lines in $\Delta^{v}$, i.e. $E^{v}\subseteq\Delta^{v}$, and (ii) that $E^{v}$ consists of a prefix of $E^{\ell}$ concatenated with a suffix of $E^{r}$, where $\ell$ and $r$ are the left and right child of $v$, respectively. See Figure 3.6.
Figure 3.6: The set of lines $E^{v}$ contributing to the lower envelope $\mathcal{E}^{v}$ are ordered by decreasing slope. Furthermore, the lower envelopes $E^{\ell}$ and $E^{r}$ of the left and right child of $v$ intersect in a single point pp. Therefore, $E^{v}$ consists of a prefix of $E^{\ell}$ concatenated with a suffix of $E^{r}$.
The main idea is then to annotate each node $v$ with this ordered set $E^{v}$. Line intersection queries can then be answered in $\mathcal{O}(\log k)$ time by using the $E^{v}$ set of the root. To insert or delete a line $\Delta_{j}$, we follow a root-to-leaf path. While walking back up the search path, we recompute $E^{v}$ from $E^{\ell}$ and $E^{r}$. In particular, by computing the intersection point of $\mathcal{E}^{\ell}$ and $\mathcal{E}^{r}$ and a constant number of split and concatenate operations on the involved ordered sets $E^{v}$. These operations can be implemented in $\mathcal{O}(\log k)$ time, thus leading to an $\mathcal{O}(\log^{2}k)$ update time.
Observe that the total size of all these $E^{v}$ sets is $\mathcal{O}(k\log k)$. Overmars and van Leeuwen reduce the space usage to $\mathcal{O}(k)$ by not explicitly storing the $E^{v}$ sets. Instead, they make sure that every node only stores the lines from $E^{v}$ that were not stored by the parent of $v$, and the relevant sets $E^{v}$ are reconstructed when performing updates. Furthermore, concatenating a prefix of $E^{\ell}$ with a suffix of $E^{r}$ to form $E^{v}$ typically destroys $E^{\ell}$ and $E^{r}$. They carefully describe how and which information to maintain to restore the sets appropriately.
Our data structure.
We make two small, but important, changes to the above data structure.
  • We observe that in our case the (line representing) $\Delta_{j}(\tau)$ never has greater slope than $\Delta_{j+1}$, because $\Delta_{j}(\tau)=\Delta^{\prime}_{j}(\tau)+\Delta_{j+1}(\tau)$ and $\Delta^{\prime}_{j}$ has a non-positive slope. This means that, at any time, the lines $\Delta_{s},\ldots,\Delta_{t}$ are already ordered by decreasing slope. Hence, we can use the indices of the functions to do the routing; i.e. each node $v$ will store the index $s^{v}$ of the rightmost leaf in its left subtree rather than its slope.
  • We extend the data structure so that given a linear function $f$ with non-positive slope and a value $j\in\{s,\ldots,t\}$, we can efficiently add $f$ to all linear functions $\Delta_{j},\ldots,\Delta_{t}$.
    Observe that we are adding $f$ to the lines that already have the smallest slope, so this does not change the overall ordering, and adding $f$ to all functions in an ordered set $\Delta^{v}$ does not change the ordering of the (slopes of the) lines inside $\Delta^{v}$ either. Moreover, the combinatorial structure of its lower envelope $\mathcal{E}^{v}$ remains unchanged: a function $\Delta_{a}$ is the $h^{\mathrm{th}}$ function in $E^{v}$ if and only if $\Delta_{a}+f$ is the $h^{\mathrm{th}}$ function in the lower envelope of $\{\Delta_{c}+f\mid\Delta_{c}\in E^{v}\}$.
    Hence, each node $v$ in our tree will store some additional linear function $f^{v}$ that we still have to add to all functions in the subtree rooted at $v$. As just argued, this allows us to represent the lower envelope $\mathcal{E}^{v}$ corresponding to node $v$ using $E^{v}$ and $f^{v}$.
    We will use the same augmentation in the binary search trees that represent $E^{v}$ themselves; each node $\nu$ will store some linear function $f^{\nu}$ that should still be added to all functions stored in its subtree. We explicitly store these sets $E^{v}$ in persistent red black trees using path copying [driscoll89makin], so that we can still have access to the original sets $E^{\ell}$ and $E^{r}$ after “combining” them into $E^{v}$. This does not affect the running times: we can split, concatenate, and search using these structures in $\mathcal{O}(\log k)$ time, but it increases the total size of our structure to $\mathcal{O}(k\log k)$.
    Using the above representation, we can still compute the intersection of a (query) line with $\mathcal{E}^{v}$ in $\mathcal{O}(\log k)$ time: such a search follows some root-to-leaf path in the tree representing $E^{v}$, so at every step we can locally apply the offset $f^{\nu}$ corresponding to the visited node $\nu$.
    Similarly, inserting or deleting a function still takes $\mathcal{O}(\log^{2}k)$ time. When we visit node $v$, we “push” its offset towards its children. When moving back up the path, we compute the intersection point between two envelopes $\mathcal{E}^{\ell}$ and $\mathcal{E}^{r}$. This involves some simultaneous root-to-leaf path traversal of traversals of the trees representing $E^{\ell}$ and $E^{r}$, so we can again locally apply the offsets involved. Hence, this still takes $\mathcal{O}(\log k)$ time. We then split and concatenate $E^{\ell}$ and $E^{r}$ into $E^{v}$ in additional $\mathcal{O}(\log k)$ time. (Note that this is where we use that every node $\nu$ in the tree representing $E^{v}$ can store its own offset $f^{\nu}$). Hence, the total time required is still $\mathcal{O}(\log^{2}k)$.
    To add $f$ to all functions $\Delta_{j},\ldots,\Delta_{t}$, we simply add $f$ to $f^{r}$, for all $\mathcal{O}(\log k)$ nodes $r$ hanging from the search path to $\Delta_{j}$ (in particular if the search path visits a node $v$ and its left child, we add $f$ to the right child $r$ of $v$). Note that we can determine which child of node $v$ to visit by comparing $j$ against the index $s^{v}$. When we walk back up the path, we recompute $E^{v}$ from $E^{\ell}$ and $E^{r}$ as before. Hence, the total time required is $\mathcal{O}(\log^{2}k)$.
Using the data structure.
We maintain the functions $\Delta_{s},\ldots,\Delta_{t}$ in the above data structure. The lower envelope $\mathcal{E}^{v}$ represented by the root $v$ is then exactly the function $g_{i}$. Hence, we can compute the root of $g_{i}$ by a line-intersection query on $\mathcal{E}^{v}$. To insert a new point $b_{j}$, we insert the linear function $\Delta_{j}$ into the structure, and increment the existing functions $\Delta_{j+1},\ldots,\Delta_{t}$ by $\Delta^{\prime}_{j}$. This takes $\mathcal{O}(\log^{2}k)$ time. Deleting a point is analogous. ∎
The main algorithm.
Our main algorithm sweeps the space of all possible translations, while maintaining an optimal matching $\phi$ for the current translation $\tau$, a representation of the current piece of the function $f$ (i.e., the linear function $f^{\prime}$ for which $f(\tau)=f^{\prime}(\tau)$), and the best translation $\tau^{*}\leq\tau$ found so far. To support the sweep, we also maintain a Lemma 3.11 data structure for each run $B_{i}$ induced by $\phi$, and a global priority queue. The Lemma 3.11 data structure allows us to efficiently obtain the next type (ii) event of a run $B_{i}$. The global priority queue stores all type (i) events, as well as the first type (ii) event of each run.
Figure 3.7: We sweep the domain of $f$, while maintaining a representation of the current piece $f^{\prime}$ of $f$, and the best translation $\tau^{*}\leq\tau$ found so far. Breakpoints correspond to type (i) or type (ii) events.
We initialize the priority queue by inserting all translations for which a pair $(b,r)\in B\times R$ coincide as type (i) events. Let $\tau_{0}$ be the first such event. For a translation $\tau<\tau_{0}$, the matching $\phi$ that assigns $b_{i}$ to $r_{i}$ is optimal (by Lemma 3.1). Hence, we use $\phi$ as the initial matching. We compute the corresponding function $f^{\prime}$ expressing the cost of $\phi$, construct the data structure of Lemma 3.11 on the single run induced by $\phi$, and query it for its first type (ii) event. We add this event to the priority queue. All of this can be done in $\mathcal{O}(mn)$ time.
To handle a type (i) event involving point $b_{j}$, we remove it from the data structure for its run and add it back in the same place with its updated linear function $\Delta_{j}^{\prime}(\tau)$. We query the data structure to find the next type (ii) event of the run $B_{i}$ containing $b_{j}$, and update the event of $B_{i}$ in the global priority queue if needed. Finally, if $b_{j}$ is aligned with $\phi(b_{j})$ in the event, we update $f^{\prime}$ by adding the function $2(b_{j}+\tau-\phi(b_{j}))$ and evaluate it. Handling an event of type (i) takes $\mathcal{O}(\log n+\log^{2}m)$ time, as it involves a constant number of operations in the global priority queue, each taking $\mathcal{O}(\log(nm))=\mathcal{O}(\log n)$ time, and a constant number of operations involving the Lemma 3.11 data structures, each taking $\mathcal{O}(\log^{2}m)$ time.
To handle a type (ii) event where the matching changes for points $b_{j},\ldots,b_{t}\in B_{i}$, we remove each point from the data structure for $B_{i}$ and then add them to the run they are now a part of (which can be either the existing run $B_{i+1}$ or a new run in between $B_{i}$ and $B_{i+1}$). This takes $\mathcal{O}(\log^{2}m)$ time per point, but as argued in Lemma 3.9 each point can only be involved in $\mathcal{O}(n)$ events of this type, so over all events, this takes $\mathcal{O}(nm\log^{2}m)$ time. We then recompute the type (ii) events corresponding to the at most two affected runs in $\mathcal{O}(\log m)$ time, and update them in the global priority queue in $\mathcal{O}(\log n)$ time. Here, we update $f^{\prime}$ by adding the (linear) cost function $\Delta_{i}(\tau)$ associated with the event.
Thus, we handle a total of $\mathcal{O}(nm)$ events of type (i), each taking $\mathcal{O}(\log n+\log^{2}m)$ time, and $\mathcal{O}(nm)$ events of type (ii), which take a total of $\mathcal{O}(nm(\log n+\log^{2}m))$ time as well.
Once we have processed all events, the algorithm has found an optimal translation $\tau^{*}$. We run the sweep once more from the start, and stop at translation $\tau^{*}$, then report the current matching $\phi$ as an optimal matching. Together with Theorem 3.3, this thus establishes Theorem $1.$
4 Lower Bound in One Dimension
In the Orthogonal Vectors problem (OV), we are given two sets of vectors $X,Y\subseteq\{0,1\}^{d}$ with $n=|X|=|Y|$ and the task is to decide whether there exist $x\in X$ and $y\in Y$ with $x\cdot y=0$, where $x\cdot y=\sum_{i=1}^{d}x[i]\cdot y[i]$. A naive algorithm solves this problem in time $\mathcal{O}(|X|^{2}d)$.
Hypothesis 4.1 (Orthogonal Vectors Hypothesis (OVH) [ov-seth, icm-survey]).
No algorithm solves the Orthogonal Vectors problem in time $\mathcal{O}(n^{2-\delta}d^{c})$ for any constants $\delta,c>0$.
In this section, we prove the following theorem.
Theorem 4.2.
Assuming OVH, for any constant $\delta>0$ there is no algorithm that, given sets $B,R\subseteq\mathbb{R}$ of size $|R|\geq|B|=\Omega(n)$, computes $\mathrm{EMDuT}(B,R)$ in time $\mathcal{O}(n^{2-\delta})$. This even holds with the additional restriction $B,R\subseteq\{0,1,\ldots,\mathcal{O}(n^{4})\}$.
Observe that this immediately implies Theorem˜1.2 because each coordinate is bounded by a polynomial. Hence, from now on, we focus on the proof of Theorem˜4.2.
We start by briefly sketching the reduction. As a building block, we design vector gadgets that for two vectors $x\in X\subseteq\{0,1\}^{d}$ and $y\in Y\subseteq\{0,1\}^{d}$ are sets of points $B(x)$ and $R(y)$ such that $\mathrm{EMDuT}(B(x),R(y))=0$ if $x$ and $y$ are orthogonal and $\mathrm{EMDuT}(B(x),R(y))\geq 1$ otherwise. This gadget is constructed by encoding $x[i]$ and $y[i]$ coordinate-wise with the geometric patterns presented in Figure˜4.1. Importantly, these gadgets share the same barycentre, therefore, when two such patterns are translated far apart (i.e., their convex hulls do not intersect), then their $\mathrm{EMD}$ does not depend on the coordinates of the vectors. In Section˜4.1 we formally construct these vector gadgets.
This allows us to construct a gadget such that for each translation $\tau$ only a single pair of vector gadgets overlaps, and the $\mathrm{EMD}$ between any other pair of vector gadgets does not depend on the coordinates of the vectors. In Section˜4.2 we show construct $B$ and $R$ as such collection of appropriately-spaced copies of gadgets $B(x)$ and $R(y)$ for $x\in X$ and $y\in Y$ in such a way that there exists a threshold $\Lambda$ such that $\mathrm{EMDuT}(B,R)=\Lambda$ if there are two orthogonal vectors in $X\times Y$ and $\mathrm{EMDuT}(B,R)\geq\Lambda+1$ otherwise. We formally prove these properties in Sections˜4.4 and 4.3.
4.1 Vector Gadgets
We construct two different types of gadgets depending on whether a vector belongs to set $X$ or $Y$ (see Figure˜4.1 for illustration):
Definition 4.3 (Red Vectors).
For any vector $x\in\{0,1\}^{d}$, define a set of points $R(x)$ to consist of:
  • $8d$ points at the coordinate 0,
  • $8d$ points at the coordinate $4d+1$, and
  • for every $i\in\{1,\ldots,d\}$:
    • if $x[i]=0$, add points $\{4i-3,4i-2,4i-1,4i\}$,
    • if $x[i]=1$, add points $\{4i-2,4i-1\}$.
Definition 4.4 (Blue Vectors).
For any vector $y\in\{0,1\}^{d}$, define a set of points $B(y)$ to consist of:
  • one point at the coordinate 0,
  • one point at the coordinate $4d+1$, and
  • for every $i\in\{1,\ldots,d\}$:
    • if $y[i]=0$, add points $\{4i-2,4i-1\}$,
    • if $y[i]=1$, add points $\{4i-3,4i\}$.
Figure 4.1: Gadgets for red and blue vectors in $d=2$. The top figure shows $R(x)$ for $x=(1,0)$, and the bottom figure shows $B(y)$ for $y=(0,1)$. Since $x$ and $y$ are orthogonal, each blue point corresponds to a red point with the same coordinate.
Next, we show that the above gadgets simulate the orthogonality.
Lemma 4.5.
Let $w=4d+1$ be the width of the gadget. Let $x,y\in\{0,1\}^{d}$ be $d$-dimensional vectors.
  1. 1.
    If $x$ and $y$ are orthogonal then $\mathrm{EMD}(B(y),R(x))=0$.
  2. 2.
    If $x$ and $y$ are not orthogonal then $\mathrm{EMD}(B(y)+\tau,R(x))\geq\max\{1,|\tau|\}$ for all $\tau\in\mathbb{R}$.
Moreover, if $|\tau|\geq w$, then we have $\mathrm{EMD}(B(y)+\tau,R(x))=|\tau|\cdot c_{1}-c_{2}$, where $c_{1}=2(d+1)$ and $c_{2}=4d^{2}+7d+1$.
Proof.
For the proof of Property 1, assume that $x$ and $y$ are orthogonal. We construct a matching such that $\mathrm{EMD}(B(y),R(x))=0$. Fix an index $i\in\{1,\ldots,d\}$. If $y[i]=1$, then $x[i]=0$; therefore, points at the coordinates $\{4i-3,4i\}$ exist in $R(x)$ and we can match the points of $B(y)$ at these coordinates. If $y[i]=0$, then note that the points at coordinates $\{4i-2,4i-1\}$ of $R(x)$ always exist. Hence, we can precisely match the points of $B(y)$ at these coordinates.
Next, we prove Property 2. Consider any $\tau\in\mathbb{R}$. At least one point of $B(y)$ (namely the leftmost or the rightmost point) must be matched by an edge of length $|\tau|$. This proves Property 2 if $|\tau|\geq 1$. So consider the case $|\tau|<1$. If $x$ and $y$ are not orthogonal, there exists an $i\in\{1,\ldots,d\}$ such that $x[i]=y[i]=1$. By construction, $B(y)$ contains points at the coordinates $4i$ and 4i34i-3, but $R(x)$ does not contain such points. Thus, the point $4i$ in $B(y)$ is matched to a point in distance at least $1-|\tau|$. In addition to the cost of $|\tau|$ incurred by the leftmost or rightmost point of $B(y)$, we obtain $\mathrm{EMD}(B(y),R(x))\geq 1$.
Finally, we focus on the proof of the last property. Consider any $\tau\in\mathbb{R}$ with $|\tau|\geq w$. Observe that $B(y)+\tau$ contains $2(d+1)$ points in total. The closest point in $R(x)$ to each of them is at the coordinate 0 or $w$. Because there are $8d$ such points, there are enough points to match each point in $B(y)$ to the closest point in $R(x)$. Hence, $\mathrm{EMD}(B(y)+\tau,R(x))$ matches each point of $B(y)$ to a point at the coordinate 0 or $w$ of $R(x)$. If $\tau\geq w$, the cost of the matching is
$\mathrm{EMD}(B(y)+\tau,R(x))$ $=2\tau-w+\sum_{i=1}^{d}\tau-w+(4i-2-y[i])+\sum_{i=1}^{d}\tau-w+(4i-1+y[i])$
$=2(d+1)\tau-4d^{2}-7d-1.$
Consequently, we can take $c_{1}=2(d+1)$ and $c_{2}=4d^{2}+7d+1$. It remains to check that the above choice is correct for $\tau\leq-w$. In that case, the cost of the matching is
$\mathrm{EMD}(B(y)+\tau,R(x))$ $=2|\tau|-w+\sum_{i=1}^{d}|\tau|-(4i-2-y[i])+\sum_{i=1}^{d}|\tau|-(4i-1+y[i])$
$=2(d+1)|\tau|-4d^{2}-7d-1.$
Thus, the same choice of $c_{1}$ and $c_{2}$ works. ∎
4.2 Reduction
Now we use the vector gadgets from the previous section to reduce the Orthogonal Vectors problem to $\mathrm{EMDuT}$. Specifically, given an OV instance $X,Y\subseteq\{0,1\}^{d}$ such that |X|=|Y|=n1|X|=|Y|=n-1, we construct sets $B,R\subseteq\mathbb{R}$ such that from $\mathrm{EMDuT}(B,R)$ we can easily infer whether $X,Y$ contains an orthogonal pair of vectors or not. Our reduction takes time $\mathcal{O}(nd)$ to construct the sets $B,R$, in particular the constructed sets have size $\mathcal{O}(nd)$. Hence, if there was an algorithm computing $\mathrm{EMDuT}(B,R)$ in time $\mathcal{O}(|R|^{2-\delta})$ for some constant $\delta>0$, then our reduction would yield an algorithm for OV running in time $\mathcal{O}((nd)^{2-\delta})$, which contradicts OVH (Hypothesis 4.1). That is, assuming OVH, $\mathrm{EMDuT}(B,R)$ cannot be computed in time $\mathcal{O}(|R|^{2-\delta})$ for any constant $\delta>0$.
For the reduction, we can assume that $n$ is even, because otherwise we can add a vector consisting exclusively of $1$s to both $X$ and $Y$. We can also assume that $d\leq n$, since otherwise the naive algorithm for OV already runs in time $\mathcal{O}(n^{2}d)=\mathcal{O}(nd^{2})$. Our reduction constructs the following point sets, for $\Delta\coloneqq 1000dn$:
  • Red Points: For the $i^{\mathrm{th}}$ vector $x_{i}\in X$, we create five red gadgets $R(x_{i})^{(1)},\ldots,R(x_{i})^{(5)}$. For each $k\in[5]$, we translate $R(x_{i})^{(k)}$ by $(i+kn)\cdot(n-1)\Delta$ and call it the $(i+kn)^{\mathrm{th}}$ red cell.
  • Blue Points: For the $j^{\mathrm{th}}$ vector $y_{j}\in Y$, we create a blue gadget $B(y_{j})$ and translate it by $j\cdot n\Delta$. This set of points is called the $j^{\mathrm{th}}$ blue cell.
We create five copies of red points for a technical reason that will become clear later (just three copies would be enough, but then we would need to argue about two types of optimal translations in the analysis). We denote the set of all red points by $R$, and the set of all blue points by $B$. This concludes the construction. Observe that $B,R$ can be constructed in time $\mathcal{O}(nd)$, as claimed, and that their coordinates are in $\{0,\ldots,\mathcal{O}(dn^{3})\}\subseteq\{0,\ldots,\mathcal{O}(n^{4})\}$. Let $c_{1}$ and $c_{2}$ be the constants (that depend on $d$) from Lemma˜4.5. Let
$\Lambda\coloneqq c_{1}\Delta\cdot n(n-2)/4+c_{2}\cdot(n-2).$
It remains to prove that the sets $X,Y$ contain orthogonal vectors if and only if $\mathrm{EMDuT}(B,R)\leq\Lambda$ (and thus from the value $\mathrm{EMDuT}(B,R)$ we can easily infer whether $X,Y$ contain orthogonal vectors).
4.3 Properties of the Construction
We now prove several useful properties of the construction described above.
Proposition 4.6.
There exist $i_{\tau},j_{\tau}\in\{0,\ldots,n-1\}$ and $\varepsilon\in(-1/2,1/2]$ such that for
$\tau\coloneqq\left((n-1)(2n+i_{\tau})-nj_{\tau}+\varepsilon\right)\cdot\Delta.$ (4.1)
we have $\mathrm{EMD}(B+\tau,R)=\mathrm{EMDuT}(B,R)$.
Proof.
Let $\lambda\coloneqq(n-1)n\Delta$ and let $\tau$ be the optimal translation. If $\tau<\lambda$, then we claim that $\mathrm{EMD}(B+\tau,R)\geq\mathrm{EMD}(B+\tau+\lambda,R)$. This holds because (i) the right endpoint of $B+\tau+\lambda$ is to the left of right endpoint of $R$, and (ii) because $R$ is periodic, with period $\lambda$, each edge of the matching $\mathrm{EMD}(B+\tau,R)$ can be shifted by $\lambda$ (or matched to a strictly closer point, if the shifted point is unavailable).
An analogous argument ensures that when $\tau>2\lambda$, then $\mathrm{EMD}(B+\tau,R)\geq\mathrm{EMD}(B+\tau-\lambda,R)$. Hence we can assume that $\tau\in[\lambda,2\lambda]$. Similar arguments show also show that for $\tau\in[\lambda,2\lambda]$ we have $\mathrm{EMD}(B+\tau,R)=\mathrm{EMD}(B+\tau+\lambda,R)$.
Observe, that for $\tau\in[\lambda,2\lambda]$, there exist $k,j\in\{0,\ldots,n-1\}$ and $\varepsilon\in(-1/2,1/2]$ such that:
$\tau=((n-1)\cdot(n+k+1)-j+\varepsilon)\Delta.$
Let $i_{1}:=k+j-n+1$ and $i_{2}:=k+j+1$ and observe that either $i_{1}$ or $i_{2}$ is in $\{0,\ldots,n-1\}$ and is a valid choice of $i$. ∎
From now, we assume that our optimal translation is of the form (4.1) and the parameters $i_{\tau},j_{\tau}$ and $\varepsilon$ are known. The idea behind this form is to have the property that the $i_{\tau}^{\mathrm{th}}$ red cell and $j_{\tau}^{\mathrm{th}}$ blue cell “nearly-align”. Before we explain this in more detail, our goal is to show that $|\varepsilon|$ is small.
To this end, we will use the following equality:
Proposition 4.7.
For every even integer $n$, any $j_{\tau}\in\{0,\ldots,n-1\}$ and $\varepsilon\in(-1/2,1/2]$ it holds that:
$\sum_{k\in\{1,\ldots,n-1\}}\min\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert k-j_{\tau}+(n-1)+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert\}=\lvert\varepsilon\rvert+n(n-2)/4.$
Proof.
We will consider the case $j_{\tau}<n/2$ as otherwise the reasoning is analogous. Then $\lvert k-j_{\tau}\rvert\leq\lvert k-j_{\tau}+n-1\rvert$ and we need to show:
$\sum_{k=1}^{n-1}\min\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert\}=\lvert\varepsilon\rvert+n(n-2)/4.$ (4.2)
The sum in (4.2) can be stratified into:
$|\varepsilon|+\sum_{k=1}^{j_{\tau}-1}\lvert k-j_{\tau}+\varepsilon\rvert+\sum_{k=j_{\tau}+1}^{n/2+j_{\tau}-1}\lvert k-j_{\tau}+\varepsilon\rvert+\sum_{k=n/2+j_{\tau}}^{n-1}\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert$.
Now, we can deduce the sign of each absolute value, so we obtain:
$|\varepsilon|+\sum_{k=1}^{j_{\tau}-1}(j_{\tau}-k-\varepsilon)+\sum_{k=j_{\tau}+1}^{n/2+j_{\tau}-1}(k-j_{\tau}+\varepsilon)+\sum_{k=n/2+j_{\tau}}^{n-1}(n-1-k+j_{\tau}-\varepsilon).$
Observe, that the number of $\varepsilon$ with positive and negative signs is equal, hence these cancel out, which yields:
$|\varepsilon|+\sum_{k=1}^{j_{\tau}-1}(j_{\tau}-k)+\sum_{k=j_{\tau}+1}^{n/2+j_{\tau}-1}(k-j_{\tau})+\sum_{k=n/2+j_{\tau}}^{n-1}(n-1-k+j_{\tau}).$
Finally, we change the summation index and conclude:
$|\varepsilon|+\sum_{\ell=1}^{j_{\tau}-1}\ell+\sum_{\ell=1}^{n/2-1}\ell+\sum_{\ell=j_{\tau}}^{n/2-1}\ell=|\varepsilon|+2\sum_{\ell=1}^{n/2-1}\ell=|\varepsilon|+n(n-2)/4.$
Here, we used the formula $\sum_{i=1}^{m}i=m(m+1)/2$ for m=n/21m=n/2-1. ∎
Property 4.8.
If $|\varepsilon|>0.1$, then $\mathrm{EMD}(B+\tau,R)>\Lambda$.
Proof.
Let us bound the length of $\mathrm{EMD}(B+\tau,R)$. For the $k^{\mathrm{th}}$ blue cell, the distance to any closest red point is at least:
$\min_{\ell\in\mathbb{Z}}\left\{\lvert k\cdot n\Delta+\tau-\ell\cdot(n-1)\Delta\rvert-2w\right\},$
because the width of every cell is $w=4d+1$. Note that the number of points in each blue cell is exactly $2(d+1)$, hence:
$\textup{LHS}\coloneqq\mathrm{EMD}(B+\tau,R)$ $\geq\sum_{k=1}^{n-1}2(d+1)\cdot\left(\min_{\ell\in\mathbb{Z}}\left\{\lvert kn\cdot\Delta+\tau-\ell(n-1)\cdot\Delta\rvert\right\}-2w\right).$
Next, we plug in the definition of $\tau:=((n-1)(2n+i_{\tau})-nj_{\tau}+\varepsilon)\Delta$:
LHS $\geq-4(d+1)w+2(d+1)\Delta\cdot\sum_{k=1}^{n-1}\min_{\ell\in\mathbb{Z}}\left\{\lvert(n-1)(2n+i_{\tau}-j_{\tau}+k-\ell)+k-j_{\tau}+\varepsilon\rvert\right\}$.
Observe that each summand is minimized when $|2n+i_{\tau}-j_{\tau}+k-\ell|\leq 1$, hence:
LHS $\geq-4(d+1)w+2(d+1)\Delta\cdot\sum_{k=1}^{n-1}\min\left\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert k-j_{\tau}+n-1+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert\right\}$.
Next, we use Proposition˜4.7:
LHS $\geq-4(d+1)w+2(d+1)\Delta\cdot\left(|\varepsilon|+n(n-2)/4\right).\$
Finally, we use $0.1\leq|\varepsilon|\leq 0.5$ and the fact that $\Lambda=c_{1}\Delta n(n-2)/4+c_{2}(n-2)$, where $c_{1}=2(d+1)$:
LHS $\geq-4(d+1)w+2(d+1)\Delta\lvert\varepsilon\rvert+\Lambda-(n-2)c_{2}>\Lambda,$
where we used $|\varepsilon|\geq 0.1$ and $\Delta\gg 4dn$. ∎
Consider a cell $k\in\{1,\ldots,n-1\}$ of $B$. We define $\textup{NearestCell}(k)$ as:
[…]
Recall that $n$ is even, and hence all the cases are covered. We show that $\textup{NearestCell}(k)$ is the index of the red cell that is closest to the $k^{\mathrm{th}}$ cell of $B$ after translation $\tau$, meaning that the optimal matching matches all points in the $k^{\mathrm{th}}$ blue cell to points in the $\textup{NearestCell}(k)^{\mathrm{th}}$ red cell:
Property 4.9.
If $|\varepsilon|<0.1$ then each point in the $k^{\mathrm{th}}$ blue cell is matched to some point in the $\textup{NearestCell}(k)^{\mathrm{th}}$ red cell.
Proof.
Observe that for any $\tau$ that satisfies Proposition˜4.6, it holds that every cell of $B+\tau$ lies between two consecutive cells of $R$. Note that the consecutive red cells are at a shorter distance than the consecutive blue cells. Hence, in between two consecutive red cells lies at most one blue cell.
Next, observe that each red cell consists of $\geq 16d$ points. Each blue cell consists of just $2(d+1)$ points. Therefore, each point cell of $B+\tau$ is matched to a point in either the next left or the next right red cell. The distance between the $\ell^{\mathrm{th}}$ red cell and the $k^{\mathrm{th}}$ blue cell is:
$\lvert kn\Delta+\tau-\ell(n-1)\Delta\rvert$.
After plugging in the value of $\tau$ this equals:
$\lvert(2n+i_{\tau}-j_{\tau}+k-\ell)\cdot(n-1)-j_{\tau}+k+\varepsilon\rvert\cdot\Delta.$
Observe that when $\lvert k-j_{\tau}\rvert<n/2$ this is minimized for $\ell=2n+i_{\tau}-j_{\tau}+k$ because $|\varepsilon|<0.1$. When $\lvert k-j_{\tau}\rvert\geq n/2$ we have two cases based on the sign of $(k-j_{\tau})$ which matches the definition of the $\textup{NearestCell}(k)$ cell because $|\varepsilon|<0.1$. ∎
4.4 Equivalence
We prove the equivalence by showing two implications separately.
Lemma 4.10.
If every pair of vectors $x\in X,y\in Y$ is not orthogonal, then $\mathrm{EMD}(B+\tau,R)>\Lambda$.
Proof.
We assume that $\tau$ is of form defined in Proposition˜4.6 as otherwise $\mathrm{EMD}(B+\tau,R)$ is large. Moreover, by ˜4.8 we can assume that $|\varepsilon|\leq 0.1$. The distance between the left endpoint of the $k^{\mathrm{th}}$ cell of $B$ and the leftmost endpoint of the nearest cell of $R$ is given by:
$\lvert(n-1)\cdot\textup{NearestCell}(k)\cdot\Delta-nk\cdot\Delta-\tau\rvert$.
Because each pair of vectors in the instance is not orthogonal, by Lemma˜4.5, we know that the smallest distance between points in each cell is at least $1$. Hence, after plugging in the definition of $\textup{NearestCell}(k)$ and by Lemma˜4.5, the contribution of points in the $k^{\mathrm{th}}$ cell of $B$ to $\mathrm{EMD}(B+\tau,R)$ is:
[…]
Observe that $\mathrm{Cost}(k)$ is at least $c_{1}\Delta\cdot\min\{\lvert j_{\tau}-k+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert,\lvert j_{\tau}-k+(n-1)+\varepsilon\rvert\}+c_{2}$ when $k\neq j_{\tau}$. Hence, the total length of $\mathrm{EMD}(B+\tau,R)$ is at least $\sum_{k=1}^{n-1}\mathrm{Cost}(k)$ which is bounded by:
$1+(n-2)c_{2}+c_{1}\Delta\Big(\sum_{k\in\{1,\ldots,n-1\}\setminus\{j_{\tau}\}}\min\{$ $\lvert j_{\tau}-k+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert$,
[…]
We use Proposition˜4.7 to conclude that:
$\sum_{k=1}^{n-1}\mathrm{Cost}(k)\geq 1+c_{2}(n-2)+c_{1}\Delta\cdot n(n-2)/4=1+\Lambda.\qed$
Lemma 4.11.
If $x_{i}\in X$ and $y_{j}\in Y$ are orthogonal, then $\mathrm{EMDuT}(B,R)\leq\Lambda$.
Proof.
Take $\tau^{\ast}=((n-1)(2n+i)-nj)\cdot\Delta$. Observe that the $(2n+i)^{\mathrm{th}}$ cell of $R$ and the $j^{\mathrm{th}}$ cell of $B$ align. Because $x_{i}$ and $y_{j}$ are orthogonal, Lemma˜4.5 guarantees that $\mathrm{EMD}(R(x_{i}),B(y_{j}))=0$. By ˜4.9, all remaining cells are matched to their closest cell. Therefore, by Lemma˜4.5 the contribution of edges with endpoints in the $k^{\mathrm{th}}$ cell of $B$ to $\mathrm{EMD}(B+\tau^{\ast},R)$ is
$\mathrm{Cost}(k)$ $=\lvert(n-1)\textup{NearestCell}(k)\cdot\Delta-nk\cdot\Delta-\tau^{\ast}\rvert\cdot c_{1}+c_{2},$
for any $k\in\{1,\ldots,n-1\}\setminus\{j\}$. By plugging in the exact values of $\textup{NearestCell}(k)$, with similar arguments as in the previous proof, we conclude that:
$\mathrm{Cost}(k)$ $=\begin{cases}0&\text{if }k=j,\\ \lvert j-k\rvert\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j\rvert<n/2,\\ (n-1-\lvert j-k\rvert)\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j\rvert\geq n/2.\end{cases}$
Observe that equivalently, we can write $\mathrm{Cost}(k)=c_{1}\Delta\cdot\min\{\lvert j-k\rvert,\lvert n-1-j+k\rvert,\lvert n-1+j-k\rvert\}+c_{2}$ when $k\neq j$. Therefore, the total length of the matching is:
$\sum_{k=1}^{n-1}\mathrm{Cost}(k)$ $=\sum_{k\in\{1,\ldots,n-1\}\setminus\{j\}}c_{2}+\Delta c_{1}\cdot\min\{\lvert j-k\rvert,\lvert n-1-j+k\rvert,\lvert n-1+j-k\rvert\}$
$=c_{2}(n-2)+\Delta c_{1}\cdot n(n-2)/4=\Lambda.$
where in the last inequality we used Proposition˜4.7. ∎
5 Lower Bounds in Higher Dimension
In this section, we prove conditional lower bounds for approximating $\mathrm{EMDuT}$ with the $L_{1}$ or $L_{\infty}$ norm. Our lower bounds assume the popular Exponential Time Hypothesis (ETH), which postulates that the 3-SAT problem on $N$ variables cannot be solved in time $2^{o(N)}$ [ImpagliazzoP01].
Theorem 5.1.
Assuming ETH, there is no algorithm that, given $\varepsilon\in(0,1)$ and $B,R\subseteq\mathbb{R}^{d}$ of size $|B|=|R|=n$, computes a $(1+\varepsilon)$-approximation of $\mathrm{EMDuT}_{1}(B,R)$ (or $\mathrm{EMDuT}_{\infty}(B,R)$) and runs in time $(\frac{n}{\varepsilon})^{o(d)}$.
We prove our lower bounds by a reduction from the $k$-Clique problem: Given a graph $G=(V,E)$ with $N$ nodes, decide whether there exist distinct nodes $v_{1},\ldots,v_{k}\in V$ such that $(v_{i},v_{j})\in E$ for all $1\leq i<j\leq k$. Here, we always assume that $k$ is constant. A naive algorithm solves the $k$-Clique problem in time $\mathcal{O}(N^{k})$. It is well known that this running time cannot be improved to $N^{o(k)}$ assuming ETH.
Theorem 5.2 ([ChenHKX06]).
Assuming ETH, the $k$-Clique problem cannot be solved in time $N^{o(k)}$.
In our lower bounds we will use the following lemma that combines gadgets $(B_{1},R_{1}),\ldots,(B_{k},R_{k})$ into a single instance $(B,R)$ whose cost is essentially the total cost of all gadgets. To prove this lemma, we simply place the gadgets sufficiently far apart.
Lemma 5.3 (Gadget Combination Lemma).
Let $1\leq p\leq\infty$. Given sets $B_{1},R_{1},\ldots,B_{k},R_{k}\subset\mathbb{R}^{d}$ of total size $n$ with $|B_{i}|\leq|R_{i}|$ for all $i\in[k]$, in time $\mathcal{O}(nd)$ we can compute sets $B,R\subset\mathbb{R}^{d}$ of total size $n$ such that
$\mathrm{EMDuT}_{p}(B,R)=\min_{\tau\in\mathbb{R}^{d}}\sum_{i=1}^{k}\mathrm{EMD}_{p}(B_{i}+\tau,R_{i}).$
Proof.
The intuition is as follows. For a sufficiently large number $U$ we construct the sets $B\coloneqq\bigcup_{i=1}^{k}B_{i}+(U\cdot i,0,\ldots,0)$ and $R\coloneqq\bigcup_{i=1}^{k}R_{i}+(U\cdot i,0,\ldots,0)$, i.e., we place the gadgets sufficiently far apart. Then one can argue that any optimal matching must match points in $B_{i}$ to points in $R_{i}$, and thus the $\mathrm{EMDuT}$ cost splits over the gadgets as claimed.
[…]
$B\coloneqq\bigcup_{i=1}^{k}B_{i}+(U\cdot i,0,\ldots,0),\qquad R\coloneqq\bigcup_{i=1}^{k}R_{i}+(U\cdot i,0,\ldots,0).$
Note that we have
$\mathrm{EMDuT}_{p}(B,R)=\min_{\tau\in\mathbb{R}^{d}}\mathrm{EMD}_{p}(B+\tau,R)\leq\min_{\tau\in\mathbb{R}^{d}}\sum_{i=1}^{k}\mathrm{EMD}_{p}(B_{i}+\tau,R_{i}),$
where the inequality follows from restricting the matching $\phi\colon(B+\tau)\to R$ to map points in $B_{i}+\tau$ to points in $R_{i}$ for all $i$.
For the opposite direction, by considering $\tau\coloneqq(0,\ldots,0)$ and considering any matching that maps points in $B_{i}+\tau$ to points in $R_{i}$ for all $i$, we observe
$\mathrm{EMDuT}_{p}(B,R)\leq|R|\cdot\Delta\leq n\Delta.$
Now consider an optimal translation $\tau^{*}$, i.e., $\tau^{*}$ realizes $\mathrm{EMDuT}_{p}(B,R)=\mathrm{EMD}_{p}(B+\tau^{*},R)$. We claim that $\|\tau^{*}\|_{p}\leq(n+2)\Delta$. Indeed, suppose for the sake of contradiction that $\|\tau^{*}\|_{p}>(n+2)\Delta$. Then any point in $B_{i}+\tau^{*}$ has distance more than $\|\tau^{*}\|_{p}-2\Delta$ to any point in $R_{i}$. It follows that if $\tau^{*}_{1}\leq 0$, then any point in $B_{1}+\tau^{*}$ has distance more than $\|\tau^{*}\|_{p}-2\Delta$ to any point in $R$, and thus $\mathrm{EMD}_{p}(B+\tau^{*},R)>\|\tau^{*}\|_{p}-2\Delta\geq n\Delta$. This contradicts $\mathrm{EMD}_{p}(B+\tau^{*},R)=\mathrm{EMDuT}_{p}(B,R)\leq n\Delta$, as shown above. Similarly, if $\tau^{*}_{1}\geq 0$, then any point in $B_{k}+\tau^{*}$ has distance more than $\|\tau^{*}\|_{p}-2\Delta$ to any point in $R$, and we again arrive at a contradiction. Hence, we have $\|\tau^{*}\|_{p}\leq(n+2)\Delta$.
Now consider an optimal matching $\phi\colon(B+\tau^{*})\to R$. If $\phi$ matches any point in $B_{i}+\tau^{*}$ to any point in $R_{j}$ for some $i\neq j$, then it incurs a cost of at least $U-2\Delta-\|\tau^{*}\|_{p}\geq U-(n+4)\Delta>n\Delta$, contradicting our upper bound $\mathrm{EMDuT}_{p}(B,R)\leq n\Delta$. Therefore, $\phi$ matches points in $B_{i}+\tau^{*}$ to points in $R_{i}$ for all $i$, and we obtain
$\mathrm{EMDuT}_{p}(B,R)=\mathrm{EMD}_{p}(B+\tau^{*},R)=\sum_{i=1}^{k}\mathrm{EMD}_{p}(B_{i}+\tau^{*},R_{i})\geq\min_{\tau\in\mathbb{R}^{d}}\sum_{i=1}^{k}\mathrm{EMD}_{p}(B_{i}+\tau,R_{i}).$
Both directions together prove the lemma. ∎
For the readers’ convenience, in Section 5.1 as a warmup we prove the lower bound for the $L_{1}$ norm in the asymmetric setting, i.e., we allow $|B|$ to be smaller than $|R|=n$. Then in Section 5.2 we strengthen this lower bound to hold even in the symmetric setting $|B|=|R|=n$. Finally, in Section 5.3 we prove the lower bound for the $L_{\infty}$ norm in the symmetric setting.
5.1 Lower Bound for $L_{1}$ Asymmetric
Figure 5.1: The left figure illustrates gadget $(B_{i,j},R_{i,j})$. The right figure illustrates gadget $(B_{i,j}^{\prime},R_{i,j}^{\prime})$.
In this section we prove Theorem 5.1 for the $L_{1}$ norm in the asymmetric setting, i.e., we relax the condition $|B|=|R|$ to $|B|\leq|R|$.
We are given a $k$-Clique instance $G=([N],E)$. We set the dimension to $d\coloneqq k$. In what follows by $p_{i,u,j,v,b}\in\mathbb{R}^{d}$ we denote the point with coordinates, for $\ell\in[d]$,
$(p_{i,u,j,v,b})_{\ell}=\begin{cases}u&\text{if }\ell=i,\\ v&\text{if }\ell=j,\\ b&\text{otherwise.}\end{cases}$
We construct the following $2{k\choose 2}$ gadgets. For any $1\leq i<j\leq k$ we construct
$B_{i,j}$ $\coloneqq\{(0,\ldots,0)\},$ $R_{i,j}$ $\coloneqq\{p_{i,u,j,v,0}\mid(u,v)\in E\},$
$B^{\prime}_{i,j}$ $\coloneqq\{(0,\ldots,0)\},$ $R^{\prime}_{i,j}$ $\coloneqq\{p_{i,u,j,v,N}\mid(u,v)\in E\}.$
The cost of these gadgets has the following properties. [^2] Recall that $\lfloor x\rceil$ denotes the closest integer to $x$, while $[x]$ denotes $\{1,\ldots,x\}$.
Lemma 5.4.
Let $1\leq i<j\leq k$. For any $\tau\in\mathbb{R}^{d}$ we have
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq(d-2)N,$
and equality holds if $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})\in E$. Moreover, for any $\tau\in\mathbb{R}^{d}$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ we have
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq(d-2)N+1.$
Proof.
Observe that
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})=$ $\min_{(u,v)\in E}\|(0,\ldots,0)+\tau-p_{i,u,j,v,0}\|_{1}$
$=\min_{(u,v)\in E}|\tau_{i}-u|+|\tau_{j}-v|+\sum_{\ell\neq i,j}|\tau_{\ell}|$
$\geq\min_{(u,v)\in E}|\tau_{i}-u|+|\tau_{j}-v|+\sum_{\ell\neq i,j}\tau_{\ell},$
where equality holds if $\tau\in[N]^{d}$. We similarly bound
$\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq\min_{(u,v)\in E}|\tau_{i}-u|+|\tau_{j}-v|+\sum_{\ell\neq i,j}N-\tau_{\ell},$
where equality holds if $\tau\in[N]^{d}$. Summing up and bounding the absolute values by 0, we obtain
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq(d-2)N.$
If $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})\in E$, then we can pick $u,v$ with $|\tau_{i}-u|+|\tau_{j}-v|=0$, and we obtain equality.
Moreover, for any $\tau\in\mathbb{R}^{d}$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$, note that since $(\tau_{i},\tau_{j})$ has $L_{\infty}$ distance at most $1/2$ to $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)$, it has $L_{\infty}$ distance at least $1/2$ to any other grid point. In particular, $(\tau_{i},\tau_{j})$ has $L_{\infty}$ distance at least $1/2$ to any $(u,v)\in E$. Since $L_{\infty}$ distance lower bounds $L_{1}$ distance, we obtain $\min_{(u,v)\in E}|\tau_{i}-u|+|\tau_{j}-v|\geq 1/2$. This yields
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})$ $\geq 2\min_{(u,v)\in E}\Big(|\tau_{i}-u|+|\tau_{j}-v|\Big)+(d-2)N$
$\geq(d-2)N+1.\qed$
We apply the Gadget Combination Lemma to the gadgets $B_{i,j},R_{i,j},B^{\prime}_{i,j},R^{\prime}_{i,j}$ for $1\leq i<j\leq d$. The value of the $\mathrm{EMDuT}_{1}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the above lemma implies the following. If $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k})\in[N]^{d}$ has a total cost of ${d\choose 2}\cdot(d-2)N\eqcolon\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Thus, each pair of gadgets contributes cost at least (d2)N(d-2)N, and at least one pair of gadgets contributes cost at least (d2)N+1(d-2)N+1, so the total cost is at least ${d\choose 2}\cdot(d-2)N+1=\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we had a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2})$, $1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N)$, and $d=k$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{3})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.
5.2 Lower Bound for $L_{1}$ Symmetric
Now we strengthen the construction to work in the symmetric setting, where the number of blue and red points is equal. To this end, we add more blue points, and for technical reasons we also need to double the number of dimensions.
We are given a $k$-Clique instance $G=([N],E)$. We set the dimension to $d\coloneqq 2k$. In what follows by $\bar{p}_{i,u,j,v,b}\in\mathbb{R}^{d}$ and $q_{i,j}\in\mathbb{R}^{d}$ we denote the points with coordinates, for $\ell\in[d]$,
[…]
We construct the following $2{k\choose 2}$ gadgets. For any $1\leq i<j\leq k$ we construct [^3] Here, we treat $B_{i,j}$ as a multi-set, containing |E|1|E|-1 times the points $q_{i,j}$. This can be avoided by adding a tiny perturbation to each copy, which makes $B_{i,j}$ a set without significantly changing any distances.
$B_{i,j}$ $\coloneqq\{(0,\ldots,0)\}\cup\{|E|-1\text{ copies of }q_{i,j}\},$ $R_{i,j}$ $\coloneqq\{\bar{p}_{i,u,j,v,0}\mid(u,v)\in E\},$
$B^{\prime}_{i,j}$ $\coloneqq\{(0,\ldots,0)\}\cup\{|E|-1\text{ copies of }-q_{i,j}\},$ $R^{\prime}_{i,j}$ $\coloneqq\{\bar{p}_{i,u,j,v,N}\mid(u,v)\in E\}.$
The cost of these gadgets has the following properties.
Lemma 5.5.
Let $1\leq i<j\leq k$. For any $\tau\in\mathbb{R}^{d}$ we have
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq((d+4)|E|-8)N,$
with equality if $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})\in E$. For any $\tau\in\mathbb{R}^{d}$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ we have
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq((d+4)|E|-8)N+1.$
Proof.
Note that in any dimension $\ell\not\in\{i,j,i+k,j+k\}$ all points in $B_{i,j}$ and $B^{\prime}_{i,j}$ have coordinate 0, and thus the contribution of these dimensions to the total cost does not depend on the matching. For $(B_{i,j},R_{i,j})$ such a dimension $\ell$ contributes $|E|\cdot|\tau_{\ell}|\geq|E|\tau_{\ell}$ to the total cost, and for $(B^{\prime}_{i,j},R^{\prime}_{i,j})$ it contributes $|E|\cdot|N-\tau_{\ell}|\geq|E|(N-\tau_{\ell})$. Summing up over both gadgets and all $\ell\not\in\{i,j,i+k,j+k\}$ yields total cost at least (d4)|E|N(d-4)|E|N, with equality if $\tau\in[N]^{d}$.
In the remainder we can focus on the dimensions $i,j,i+k,j+k$. Projected onto these dimensions, the $L_{1}$ distance from $q_{i,j}+\tau$ to $\bar{p}_{i,u,j,v,0}$ is
$|N+\tau_{i}-u|+|N+\tau_{j}-v|+|N-\tau_{i+k}+u|+|N-\tau_{j+k}+v|\geq 4N+\tau_{i}+\tau_{j}-\tau_{i+k}-\tau_{j+k},$
where we used $|x|\geq x$. Furthermore, equality holds if $\tau\in[N]^{d}$. Analogously, one can show that projected onto the dimensions $i,j,i+k,j+k$ the $L_{1}$ distance from $q_{i,j}+\tau$ to $\bar{p}_{i,u,j,v,N}$ is
$\geq 4N-\tau_{i}-\tau_{j}+\tau_{i+k}+\tau_{j+k},$
with equality if $\tau\in[N]^{d}$. Similarly, projected onto the dimensions $i,j,i+k,j+k$ the $L_{1}$ distance from $(0,\ldots,0)+\tau$ to $\bar{p}_{i,u,j,v,0}$ (or to $\bar{p}_{i,u,j,v,N}$) is
$|\tau_{i}-u|+|\tau_{j}-v|+|\tau_{i+k}-u|+|\tau_{j+k}-v|.$
By summing up everything, and noting that the terms $\tau_{i}+\tau_{j}-\tau_{i+k}-\tau_{j+k}$ and $\tau_{i}-\tau_{j}+\tau_{i+k}+\tau_{j+k}$ cancel, we obtain
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})$
$\geq 2\min_{(u,v)\in E}|\tau_{i}-u|+|\tau_{j}-v|+|\tau_{i+k}-u|+|\tau_{j+k}-v|+(|E|-1)\cdot 2\cdot 4N+(d-4)|E|N,$
with equality if $\tau\in[N]^{d}$. Bounding the absolute values by 0 and simplifying, we obtain
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq((d+4)|E|-8)N,$
with equality if $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})\in E$.
Moreover, for any $\tau\in\mathbb{R}^{d}$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$, the point $(\tau_{i},\tau_{j})$ has $L_{\infty}$ distance at least $1/2$ to any $(u,v)\in E$, and thus $\min_{(u,v)\in E}|\tau_{i}-u|+|\tau_{j}-v|\geq 1/2$. In this case we obtain
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq((d+4)|E|-8)N+2\cdot\tfrac{1}{2}.\qed$
We apply the Gadget Combination Lemma to the gadgets $B_{i,j},R_{i,j},B^{\prime}_{i,j},R^{\prime}_{i,j}$ for $1\leq i<j\leq k$. The value of the $\mathrm{EMDuT}_{1}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the above lemma implies the following. If $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k},v_{1},\ldots,v_{k})\in[N]^{d}$ has a total cost of ${k\choose 2}\cdot((d+4)|E|-8)N=:\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Thus, each pair of gadgets contributes cost at least ((d+4)|E|8)N((d+4)|E|-8)N, and at least one pair of gadgets contributes cost at least ((d+4)|E|8)N+1((d+4)|E|-8)N+1, so the total cost is at least ${k\choose 2}\cdot((d+4)|E|-8)N+1=\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we had a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2}),1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N^{3})$, and $d=\mathcal{O}(k)$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{5})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.
5.3 Lower Bound for $L_{\infty}$ Symmetric
In this section we prove Theorem 5.1 for the $L_{\infty}$ norm, thus finishing the proof of this theorem.
We are given a $k$-Clique instance $G=([N],E)$. We set the dimension to $d\coloneqq 2k+1$. In what follows, by $\hat{p}_{i,u,j,v,b}\in\mathbb{R}^{d}$, $q_{i}\in\mathbb{R}^{d}$, and $b_{i,j}\in\mathbb{R}^{d}$ we denote the points with coordinates, for $\ell\in[d]$,
$(\hat{p}_{i,u,j,v,b})_{\ell}$ $=\begin{cases}u&\text{if }\ell\in\{i,i+k\},\\ v&\text{if }\ell\in\{j,j+k\},\\ b&\text{otherwise,}\end{cases}\qquad(b_{i,j})_{\ell}=\begin{cases}10N&\text{if }\ell\in\{i,j\},\\ -10N&\text{if }\ell\in\{i+k,j+k\},\\ 0&\text{otherwise,}\end{cases}$
$(q_{i})_{\ell}$ $=\begin{cases}10N&\text{if }\ell\in\{i,i+k\},\\ 0&\text{otherwise.}\end{cases}$
We construct the following $4(k-1)k+2{k\choose 2}$ gadgets. For any $i\in[k],s\in[2(k-1)]$ we construct
$B_{i,s}$ $\coloneqq\{(0,\ldots,0)\},$ $R_{i,s}$ $\coloneqq\{q_{i}\},$
$B^{\prime}_{i,s}$ $\coloneqq\{(0,\ldots,0)\},$ $R^{\prime}_{i,s}$ $\coloneqq\{-q_{i}\}.$
Moreover, for any $1\leq i<j\leq k$ we construct
$\hat{B}_{i,j}$ $\coloneqq\{b_{i,j}\}\cup\{|E|-1\text{ copies of }(0,\ldots,0,10N)\},$ $\hat{R}_{i,j}$ $\coloneqq\{\hat{p}_{i,u,j,v,0}\mid(u,v)\in E\},$
$\hat{B}^{\prime}_{i,j}$ $\coloneqq\{b_{i,j}\}\cup\{|E|-1\text{ copies of }(0,\ldots,0,-10N)\},$ $\hat{R}^{\prime}_{i,j}$ $\coloneqq\hat{R}_{i,j}.$
The following two lemmas analyze the properties of these gadgets.
Lemma 5.6.
Let $i\in[k]$ and $s\in[2(k-1)]$. For any $\tau\in\mathbb{R}^{d}$ we have
$\mathrm{EMD}_{\infty}(B_{i,s}+\tau,R_{i,s})+\mathrm{EMD}_{\infty}(B^{\prime}_{i,s}+\tau,R^{\prime}_{i,s})\geq 20N+|\tau_{i}-\tau_{i+k}|,$
with equality if $\tau\in[N]^{d}$.
Proof.
We have
$\mathrm{EMD}_{\infty}(B_{i,s}+\tau,R_{i,s})=\|(0,\ldots,0)+\tau-q_{i}\|_{\infty}\geq\max\{10N-\tau_{i},10N-\tau_{i+k}\},$
with equality if $\tau\in[N]^{d}$, as then the coordinates involving $10N$ dominate. Similarly, we have
$\mathrm{EMD}_{\infty}(B^{\prime}_{i,s}+\tau,R^{\prime}_{i,s})\geq\max\{10N+\tau_{i},10N+\tau_{i+k}\},$
with equality if $\tau\in[N]^{d}$. We bound their sum by
$\mathrm{EMD}_{\infty}(B_{i,s}+\tau,R_{i,s})+\mathrm{EMD}_{\infty}(B^{\prime}_{i,s}+\tau,R^{\prime}_{i,s})$
$\geq\max\{(10N-\tau_{i})+(10N+\tau_{i+k}),(10N-\tau_{i+k})+(10N+\tau_{i})\}=20N+|\tau_{i}-\tau_{i+k}|.$
We observe that equality holds if $\tau\in[N]^{d}$, as then we have
$\mathrm{EMD}_{\infty}(B_{i,s}+\tau,R_{i,s})+\mathrm{EMD}_{\infty}(B^{\prime}_{i,s}+\tau,R^{\prime}_{i,s})$
$=\max\{(10N-\tau_{i})+(10N+\tau_{i+k}),(10N-\tau_{i+k})+(10N+\tau_{i}),$
$\qquad\quad\;\;\;(10N-\tau_{i})+(10N+\tau_{i}),(10N-\tau_{i+k})+(10N+\tau_{i+k})\}=20N+|\tau_{i}-\tau_{i+k}|.\qed$
Lemma 5.7.
Let $1\leq i<j\leq k$. For any $\tau\in\mathbb{R}^{d}$ we have
$\mathrm{EMD}_{\infty}(\hat{B}_{i,j}+\tau,\hat{R}_{i,j})+\mathrm{EMD}_{\infty}(\hat{B}^{\prime}_{i,j}+\tau,\hat{R}^{\prime}_{i,j})\geq 20N|E|-2|\tau_{i}-\tau_{i+k}|-2|\tau_{j}-\tau_{j+k}|,$
with equality if $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})\in E$. For any $\tau\in\mathbb{R}^{d}$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ we have
$\mathrm{EMD}_{\infty}(\hat{B}_{i,j}+\tau,\hat{R}_{i,j})+\mathrm{EMD}_{\infty}(\hat{B}^{\prime}_{i,j}+\tau,\hat{R}^{\prime}_{i,j})\geq 20N|E|-2|\tau_{i}-\tau_{i+k}|-2|\tau_{j}-\tau_{j+k}|+1.$
Proof.
Since the last coordinate of any point in $\hat{R}_{i,j}$ (or $\hat{R}^{\prime}_{i,j}$) is 0, its distance to $(0,\ldots,0,10N)+\tau$ is at least $10N+\tau_{d}$, with equality if $\tau\in[N]^{d}$. Similarly, its distance to $(0,\ldots,0,-10N)+\tau$ is at least $10N-\tau_{d}$, with equality if $\tau\in[N]^{d}$. Therefore, all copies of $(0,\ldots,0,10N)$ and $(0,\ldots,0,-10N)$ in total contribute a cost of at least $2(|E|-1)\cdot 10N$, with equality if $\tau\in[N]^{d}$, no matter what points in $\hat{R}_{i,j}$ (or $\hat{R}^{\prime}_{i,j}$, resp.) they are assigned to.
The remaining point $b_{i,j}+\tau$ has distance to $\hat{R}_{i,j}$ (and to $\hat{R}^{\prime}_{i,j}$) of
$\geq\min_{(u,v)\in E}\max\{10N+\tau_{i}-u,10N+\tau_{j}-v,10N-\tau_{i+k}+u,10N-\tau_{j+k}+v\}.$ (5.1)
Equality holds for $\tau\in[N]^{d}$, as then then coordinates involving $10N$ dominate. We further bound (5.1) from below by
$=10N+\min_{(u,v)\in E}\max\{\tau_{i}-u,\tau_{j}-v,u-\tau_{i+k},v-\tau_{j+k}\}$
$\geq 10N+\min_{(u,v)\in E}\max\{\tau_{i}-u,u-\tau_{i},\tau_{j}-v,v-\tau_{j}\}-|\tau_{i}-\tau_{i+k}|-|\tau_{j}-\tau_{j+k}|$
$=10N+\min_{(u,v)\in E}\max\{|\tau_{i}-u|,|\tau_{j}-v|\}-|\tau_{i}-\tau_{i+k}|-|\tau_{j}-\tau_{j+k}|.$
Equality holds if $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})$.
Summing up these costs, in total we obtain
$\mathrm{EMD}_{\infty}(\hat{B}_{i,j}+\tau,\hat{R}_{i,j})+\mathrm{EMD}_{\infty}(\hat{B}^{\prime}_{i,j}+\tau,\hat{R}^{\prime}_{i,j})$
$\geq 20N|E|+2\min_{(u,v)\in E}\Big(\max\{|\tau_{i}-u|,|\tau_{j}-v|\}-|\tau_{i}-\tau_{i+k}|-|\tau_{j}-\tau_{j+k}|\Big),$
with equality if $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})$. By bounding the $\max$ term by 0, we obtain
$\mathrm{EMD}_{\infty}(\hat{B}_{i,j}+\tau,\hat{R}_{i,j})+\mathrm{EMD}_{\infty}(\hat{B}^{\prime}_{i,j}+\tau,\hat{R}^{\prime}_{i,j})\geq 20N|E|-2|\tau_{i}-\tau_{i+k}|-2|\tau_{j}-\tau_{j+k}|.$
Equality holds if $\tau\in[N]^{d}$ and $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})\in E$.
Moreover, for any $\tau\in\mathbb{R}^{d}$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$, the point $(\tau_{i},\tau_{j})$ has $L_{\infty}$ distance at least $1/2$ from any $(u,v)\in E$, and thus $\min_{(u,v)\in E}\max\{|\tau_{i}-u|,|\tau_{j}-v|\}\geq 1/2$. In this case, we obtain cost at least
$20N|E|+2\cdot\tfrac{1}{2}-2|\tau_{i}-\tau_{i+k}|-2|\tau_{j}-\tau_{j+k}|.\qed$
We now apply the Gadget Combination Lemma to the gadgets $B_{i,s},R_{i,s},B^{\prime}_{i,s},R^{\prime}_{i,s}$ for $i\in[k]$ and $s\in[2(k-1)]$ and $\hat{B}_{i,j},\hat{R}_{i,j},\hat{B}^{\prime}_{i,j},\hat{R}^{\prime}_{i,j}$ for $1\leq i<j\leq k$. The value of $\mathrm{EMDuT}_{\infty}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the cost can be read off from the above lemmas. Note that for any $i\in[k]$ the terms $|\tau_{i}-\tau_{i+k}|$ cancel, as there are 2(k1)2(k-1) gadget pairs contributing $|\tau_{i}-\tau_{i+k}|$ (one gadget pair for each $s\in[2(k-1)]$), and k1k-1 gadget pairs contributing $2|\tau_{i}-\tau_{i+k}|$ (one gadget pair for each $j\in[k],j\neq i$). The cost thus simplifies to
$\mathrm{EMD}_{\infty}(B+\tau,R)\geq 20N\cdot k\cdot 2(k-1)+20N|E|\cdot{k\choose 2}\eqcolon\Lambda.$
We obtain equality for any translation $\tau\in[N]^{d}$ with $(\tau_{i},\tau_{j})=(\tau_{i+k},\tau_{j+k})\in E$ for all $1\leq i<j\leq k$. Therefore, if $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k},v_{1},\ldots,v_{k},0)\in[N]^{d}$ has a total cost of $\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Then according to Lemma 5.7 at least one summand has an additional +1+1, so the total cost is at least $\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{\infty}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we had a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{\infty}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2}),1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N^{3})$, and $d=\mathcal{O}(k)$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{5})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.
6 Algorithms in Higher Dimensions
Given two sets $R$ and $B$ of $n$ points in the plane, Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] show how to compute a translation $\tau^{*}$ minimizing $\mathrm{EMDuT}_{1}(B,R)$ with respect to the $L_{1}$-distance in $\mathcal{O}(n^{6}\log^{3}n)$ time. We observe that their result can be generalized to point sets in arbitrary dimension $d$, leading to an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm.
Furthermore, we show that our approach can be used to obtain an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm for finding a translation that minimizes $\mathrm{EMDuT}_{\infty}(B,R)$, i.e. the Earth Mover’s Distance with respect to the $L_{\infty}$ distance. For point sets in $\mathbb{R}^{2}$, this immediately follows by “rotating the plane by $45^{\circ}$” and using the algorithm for $L_{1}$. For higher dimensions this trick is no longer immediately applicable. However, we show that our algorithm can also directly be applied to the $L_{\infty}$ distance, even for point sets in $\mathbb{R}^{d}$ with $d>2$.
Earth Mover’s Distance without Translation.
We first describe an algorithm to compute $\mathrm{EMD}_{p}(B,R)$ in $\mathbb{R}^{d}$. Note that we assume to work in the Real RAM model, hence we need a strongly-polynomial algorithm. Naively, one can achieve that in $\mathcal{O}(m^{2}n)$ time by computing the bipartite graph, and solving maximum weight matching in bipartite graph in strongly polynomial time by Edmonds and Karp [edmonds1972theoretical]. Here, however, we can use the fact that points are in $\mathbb{R}^{d}$. To the best of our knowledge, the best algorithm in this setting is due to Vaidya [vaidya89geomet_helps_match]. However, he only considers the case when both point sets are in $\mathbb{R}^{2}$ and have size $n=m$ in $\mathbb{R}^{2}$. He shows that one can compute $\mathrm{EMD}_{p}(B,R)$ (with $p\in\{1,\infty\}$) in $\mathcal{O}(n^{2}\log^{3}n)$ time in this setting. Furthermore, he states (without proof) that for point sets in $\mathbb{R}^{d}$, that the running time increases by at most $\mathcal{O}(\log^{d}n)$. Next, we briefly sketch the algorithm and fill in the missing details for the higher-dimensional setting, to obtain the following result:
Theorem 6.1.
Given a set $B$ of $m$ points in $\mathbb{R}^{d}$, and a set of $n\geq m$ red points in $\mathbb{R}^{d}$, there is an $\mathcal{O}(n^{2}\log^{d+2}n)$ time algorithm to compute $\mathrm{EMD}_{p}(B,R)$, for $p\in\{1,\infty\}$.
Proof.
Vaidya’s algorithm is a particular implementation of the Hungarian Method [kuhn1955hungarian]. Hence, to apply it we introduce nmn-m additional “dummy” points in $B$. We define the distance from a dummy point to any other point in $R$ to be zero. We stress that the dummy points are only present in the graph representation, and are not physically in $\mathbb{R}^{d}$. The main algorithm proceeds in phases, in each of which the current matching grows by one new pair. Hence, the algorithm completes after $n$ phases. In each phase, each point $q\in B\cup R$ is assigned a weight $w_{q}$ (the current value of the point in the dual LP-formulation of the problem). The algorithm then maintains a subset $R^{\prime}$ of red points (initially $R^{\prime}=R$), and a subset $B^{\prime}$ of blue points (initially, $B^{\prime}$ is the set of unmatched blue points), and repeatedly computes the bichromatic closest pair $(b^{*},r^{*})$ with respect to the weighted distance function $d(b,r)=\|b-r\|_{p}-w_{b}-w_{r}$. This pair is either added to the matching (if $r^{*}$ is also unmatched), or the weights of $b^{*}$ and $r^{*}$ are updated, $r^{*}$ is removed from $R^{\prime}$, and $b^{*}$ is added to $B^{\prime}$. Since there are only $n$ points in one set, a phase consists of at most $\mathcal{O}(n)$ such steps. Hence, given a dynamic data structure storing $R^{\prime}$ and $B^{\prime}$ that
  • maintains the bichromatic (weighted) closest pair among $R^{\prime}\cup B^{\prime}$,
  • can be built in $P(n)$ time (and thus uses at most $P(n)$ space),
  • an insertion of a blue point into $B^{\prime}$ in (amortized) $I(n)$ time, and
  • a deletion of a red point from $R^{\prime}$ in (amortized) $D(n)$ time,
the algorithm runs in $\mathcal{O}(n(P(n)+n(I(n)+D(n))))$ time. As we argue next, for (weighted) points in $\mathbb{R}^{d}$, there is such a data structure with $P(n)=\mathcal{O}(n\log^{d}n)$, and $I(n)\leq D(n)=\mathcal{O}(\log^{d+2}n)$. The theorem then follows.
We describe a dynamic data structure for weighted nearest neighbor (NN) queries and then apply a recent result of Chan [chan20dynam_gener_closes_pair] to turn this into a dynamic bichromatic closest pair data structure.
Let $P^{\prime}$ be a set of weighted points in $\mathbb{R}^{d}$, let $q$ be a weighted point, and consider the positive orthant $A(q)=\{a\mid q_{i}\leq a_{i}\text{ for all }i\in\{1,\ldots,d\}\}\subset\mathbb{R}^{d}$ with respect to $q$ (i.e. all points dominating $q$). Let $w^{\prime}_{z}=-w_{p^{\prime}}+\sum_{i=1}^{d}p^{\prime}_{i}$ (see Figure 6.1). Now observe that the point $p^{*}\in P^{\prime}\cap A(q)$ minimizing $\|p^{\prime}-q\|-w_{p^{\prime}}-w_{q}$ is the point in $A(q)$ with minimum $w^{\prime}_{p^{\prime}}$ value.
Figure 6.1: An example of a $L_{1}$ unit ball centered at $q$ in $\mathbb{R}^{2}$. Observe that all points on the line through $p^{\prime}$ with slope minus one have the same value $\sum_{i=1}p^{\prime}_{i}$. For a weighted point $p^{\prime}\in P^{\prime}$ that lies in $A(q)$ (the yellow region), the weighted $L_{1}$-distance between $q$ and $p^{\prime}$ is given by $w^{\prime}_{p}-w^{\prime}_{q}$, hence the point from $P^{\prime}\cap A(q)$ with minimum $w^{\prime}_{p}$ value is the nearest neighbor of $q$.
So, we store the points in $P^{\prime}$ in a $d$-dimension range tree [preparata2012computational], in which every subtree is annotated with the point with the minimum $w^{\prime}_{p}$ value among its descendants. We can query the range tree for the nearest point among $P^{\prime}\cap A(q)$ with respect to the weighted $L_{1}$ distance in $\mathcal{O}(\log^{d}n)$ time. By implementing the trees using generalized balanced trees [andersson99gener_balan_trees] we can support insertions and deletions in amortized $\mathcal{O}(\log^{d}n)$ time as well. We use an analogous approach for the other $2^{d}-1$ orthants around $q$, thus allowing us to answer weighted NN queries with respect to the $L_{1}$ distance in the same time as above. We can handle the (weighted) $L_{\infty}$ distance analogously by defining appropriate “orthants” around $q$. If $P^{\prime}$ contains any dummy points, we store them separately in a binary search tree, ordered by weight. The minimum weight dummy point is an additional candidate nearest neighbor for $q$, and we return the overall closest point to $q$. When $q$ itself would be a dummy point (and thus has no real location), we simply obtain the minimum weight point among $P^{\prime}$ as the nearest point. Hence, it follows we can build the structure in $P_{0}(n)=O(n\log^{d}n)$ time, query the nearest neighbor in $Q_{0}(n)=O(\log^{d}n)$ time, and delete (and insert) points in $D_{0}(n)=O(\log^{d}n)$ time.
Chan [chan20dynam_gener_closes_pair] shows how to turn a dynamic nearest neighbor searching data structure with preprocessing time $P_{0}(n)$, query time $Q_{0}(n)$, and deletion time $D_{0}(n)$ into a fully dynamic bichromatic closest pair data structure. In our particular case we obtain
  • construction time $P(n)=\mathcal{O}(nQ_{0}(n)+P_{0}(n))=\mathcal{O}(n\log^{d}n)$,
  • amortized insertion time $I(n)=\mathcal{O}(Q_{0}(n)\log n+(P_{0}(n)/n)\log n)=\mathcal{O}(\log^{d+1}n)$, and
  • amortized deletion time $D(n)=\mathcal{O}(Q_{0}(n)\log^{2}n+(P_{0}(n)/n)\log^{2}n+D_{0}(n)\log n)=\mathcal{O}(\log^{d+2}n)$.
Plugging this into Vaidia’s algorithm we thus obtain an $\mathcal{O}(n^{2}\log^{d+2}n)$ time algorithm to compute $\mathrm{EMD}(B,R)$. ∎
Earth Mover’s Distance under Translation in $L_{1}$.
The sets $B$ and $R$ are aligned in dimension $i$, or $i$-aligned for short, if there is a pair of points $b\in B,r\in R$ for which $b_{i}=r_{i}$. Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] show that for two point sets in $\mathbb{R}^{2}$, there exists an optimal translation $\tau^{*}$ that aligns $B$ and $R$ in both dimensions. They explicitly consider all $\mathcal{O}((nm)^{2})$ translations that both $1$-align and $2$-align $B+\tau$ and $R$. For each such a translation $\tau$, computing an optimal matching can then be done in $\mathcal{O}(n^{2}\log^{3}n)$ time [vaidya89geomet_helps_match], thus leading to an $\mathcal{O}(n^{4}m^{2}\log^{3}n)$ time algorithm. We now argue that we can generalize the above result to higher dimensions.
Theorem 6.2.
Given $B$ and $R$ we can find an optimal translation $\tau^{*}$ realizing $\mathrm{EMDuT}_{1}(B,R)$ in $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time.
Proof.
Recall the definition of the cost function
$\mathcal{D}_{B,R,1}(\phi,\tau)=\sum_{b\in B}L_{1}(b+\tau,\phi(b))=\sum_{b\in B}\sum_{i=1}^{d}|b_{i}+\tau_{i}-\phi(b)_{i}|.$
For a fixed matching $\phi$, this is a piecewise linear function in $\tau$. In particular, $\mathcal{D}_{B,R,1}(\phi,\tau)$ is a sum of piecewise linear functions $f_{b,i}(\tau)=|b_{i}+\tau_{i}-\phi(b)_{i}|$. For each such a function there is a hyperplane $h_{b,\phi(b),i}$ in $\mathbb{R}^{d}$ given by the equation $\tau_{i}+b_{i}-\phi(b)_{i}=0$, so that for a point (translation) $\tau\in\mathbb{R}^{d}$ on one side of (or on) the hyperplane, $f_{b,i}(\tau)$ is linear in $\tau$ (i.e. on one side we have $f(\tau)=\tau_{i}+b_{i}-\phi(b)_{i}$, whereas on the other side we have $f(\tau)=-\tau_{i}-b_{i}+\phi(b)_{i}$). Let $H_{\phi}=\{h_{b,\phi(b),i}\mid b\in B,i\in\{1,\ldots,d\}\}$ denote the set of all such hyperplanes, and consider the arrangement $\mathcal{A}(H_{\phi})$. It follows that in each cell of $\mathcal{A}(H_{\phi})$, the function $\mathcal{D}_{B,R,1}(\phi,\tau)$ is a linear function in $\tau$, and that $\mathcal{D}_{B,R,1}(\phi,\tau)$ thus has its minimum at a vertex of $\mathcal{A}(H_{\phi})$.
We extend the set of hyperplanes $H_{\phi}$ to include the hyperplane $h_{b,r,i}$ for every pair $(b,r)\in B\times R$, and every $i\in\{1,\ldots,d\}$, rather than just the pairs $(b,\phi(b))$. Let $H$ be the resulting set. A minimum of $\mathcal{D}_{B,R,1}(\phi,\tau)$ still occurs at a vertex of $\mathcal{A}(H)$ (as $\mathcal{A}(H)$ includes all vertices of $\mathcal{A}(H_{\phi})$). Moreover, observe that $H$ now actually contains the hyperplanes $H_{\phi}$, for every matching $\phi\in\Phi$, so also those of an optimal matching $\phi^{*}$. It thus follows that such a global minimum $\mathcal{D}_{1}(\phi^{*},\tau^{*})$ occurs at a vertex $\tau^{*}$ of $\mathcal{A}(H)$.
So, to compute an optimal matching $\phi^{*}$ and its $\tau^{*}$ (and thus $\mathrm{EMDuT}(B,R)$) we can
  1. 1.
    explicitly compute (all vertices of) $\mathcal{A}(H)$,
  2. 2.
    for each such a vertex $\tau\in\mathcal{A}(H)$ (which is some candidate translation), compute an optimal matching $\phi_{\tau}$ between the sets $B+\tau$ and $R$, and
  3. 3.
    report the matching (and corresponding translation) that minimizes total cost.
The set $H$ contains mndmnd hyperplanes, and thus $\mathcal{A}(H)$ contains $\mathcal{O}((mnd)^{d})=\mathcal{O}(m^{d}n^{d})$ vertices. Computing $\mathcal{A}(H)$ takes $\mathcal{O}(m^{d}n^{d})$ time [bkos-cgaa-08, edelsbrunner86const_arran_lines_hyper_applic]. For each such a vertex (translation), we can compute an optimal matching in $\mathcal{O}(n^{2}\log^{d+2}n))$ time using the algorithm from Theorem 6.1. This thus yields an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm in total. ∎
EMD under translation in $L_{\infty}$.
We present the following analog for $L_{\infty}$.
Theorem 6.3.
Given $B$ and $R$ we can find an optimal translation $\tau^{*}$ realizing $\mathrm{EMDuT}_{\infty}(B,R)$ in $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time.
Proof.
We use a similar approach as in Theorem 6.2; i.e. we prove that there is a set $H$ of $\mathcal{O}(mnd^{2})$ hyperplanes in $\mathbb{R}^{d}$, so that for any matching $\phi$, there is a minimum cost translation that is a vertex of the arrangement $\mathcal{A}(H)$. We can thus again compute such an optimal matching (and the translation) by trying all $\mathcal{O}(m^{d}n^{d})$ vertices. This yields an $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time algorithm as claimed. What remains is to describe the set of hyperplanes $H$.
Fix a matching $\phi$. We have
$\mathcal{D}_{B,R,\infty}(\phi,\tau)=\sum_{b\in B}L_{\infty}(b+\tau,\phi(b))=\sum_{b\in B}\max_{i=1}^{d}|b_{i}+\tau_{i}-\phi(b)_{i}|,$
which is again a piecewise linear function in $\tau$, as it is a sum of piecewise linear functions. Each individual function is now of the form $f_{b,r}=\max_{i=1}^{d}|b_{i}+\tau_{i}-r_{i}|$, where $b\in B$, and $r=\phi(b)$. For each such a function, there are now at most $\mathcal{O}(d^{2})$ hyperplanes that subdivide $\mathbb{R}^{d}$ into regions in which $f_{b,r}$ is given by a single linear function. In particular, the $d$ hyperplanes given by $\tau_{i}-b_{i}+r_{i}=0$, for any $i\in\{1,\ldots,d\}$, and $\mathcal{O}(d^{2})$ hyperplanes that we get from solving $|b_{i}+\tau_{i}-r_{i}|=|b_{j}+\tau_{j}-r_{j}|$ for each $j\neq i$ (for example $\tau_{i}-\tau_{j}+b_{i}-b_{j}-r_{i}+r_{j}=0$). Let $H$ denote the resulting set of $\mathcal{O}(mnd^{2})$ hyperplanes over all points $b\in B$, $r\in R$, and all $i\in\{1,\ldots,d\}$. It again follows that for any matching $\phi$, there is a vertex of $\mathcal{A}(H)$ that corresponds to a translation $\tau_{\phi}$ that minimizes $\mathcal{D}_{B,R,\infty}(\phi,\tau)$. Since this holds for an optimal matching $\phi^{*}$, we can thus compute $\mathrm{EMDuT}(B,R)$ in $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time. ∎